### v1

# JigsawComm: Joint Semantic Feature Encoding and Transmission for Communication-Efficient Cooperative Perception

Abstract
[…]
1 Introduction
Multi-agent cooperative perception (CP) enables connected road agents, such as connected and autonomous vehicles (CAVs) and road-side units (RSUs), to improve their understanding of the environment [cp_benefit]. By leveraging sensor data exchanged through Vehicle-to-Everything (V2X) communication [cpm], participating agents can detect objects beyond their individual sensing ranges, minimize blind spots, and perceive occluded obstacles [xu2022opv2v]. With its great potential for facilitating more informed driving decisions and enhancing road safety, CP is a critical technology for 50–90% of the vehicle market that CAVs are anticipated to comprise by 2040 [Chatziioannou2024].
Figure 1: A sample scenario from OPV2V dataset (left). Under the same bandwidth limit, prior work [hu2022where2comm] (mid) does not utilize bandwidth efficiently, leading to missed detection and reduced accuracy. JigsawComm (right) maximizes every bit’s contribution, enabling complete and accurate detection. Quantitative evaluation and discussion on the ‘cost’ of redundancy can be found in Sec. 4 and Sec. 5, respectively.
Despite this promise, CP faces a formidable challenge in communication efficiency. The available V2X bandwidth is severely limited (up to $20$ MHz for both DSRC [FCC_DSRC] and C-V2X [CV2X_ETSI_EN_303_613]), while the volume of sensor data generated is prohibitively large (e.g. 4 MB per LiDAR frame [f_cooper]). Exchanging raw sensor data is therefore infeasible while satisfying the stringent real-time perception requirements of autonomous driving (e.g. <<100 ms [ADConstraints]). As previous research also demonstrated that exchanging local detection results alone is suboptimal [xu2022opv2v], the field has largely converged on the intermediate fusion framework—exchanging and fusing Bird’s-Eye-View (BEV) features—as the most effective paradigm [xu2022opv2v]. Even so, transmitting dense deep features from multiple agents still exhausts V2X bandwidth.
The key challenge stems from the high degree of data redundancy present in these shared features, which can be classified into two forms: (i) Intra-agent redundancy, where features waste bits on semantically irrelevant information; and (ii) Inter-agent redundancy, where multiple agents transmit overlapping and unnecessary information. As shown in Fig. 1, under limited bandwidth, such inefficient communication not only leads to reduced detection coverage but can even degrade perception accuracy when lower quality data dilutes the utility of the better ones after fusion.
Existing work on communication-efficient CP fails to solve this. They typically adopt one of two strategies: (i) Compression methods, such as encoder-decoders [xu2022opv2v] or codebooks [hu2024communication, zhao2025quantv2x]. Such methods either only reduce the data size but still transmit the entire feature map, failing to address redundancy, or overly restrict the representation power to offline-trained dictionaries, undermining generalizability; (ii) Message selection methods [liu2020when2com, hu2022where2comm, hu2024communication, xu2025cosdh], choosing ‘critical’ regions to transmit but relying on simple local heuristics (e.g. local detection confidence) that are poor proxies for the final cooperative task. Crucially, both approaches ignore cross-agent redundancy. As a result, they still require significant bandwidth and are not scalable: their communication cost grows linearly with the number of agents. The problem of designing a truly communication-efficient, scalable, and accurate CP system remains open.
To design such a CP system, we argue that it must maximize the contribution of every transmitted bit towards the final perception goal. This requires that only essential, semantically meaningful, and non-redundant information is transmitted. Therefore, the feature encoding (local compression) and transmission (message selection) algorithms must be jointly designed to realize this potential. This joint design is non-trivial. Removing semantic, cross-agent redundancy is a challenging ‘chicken-and-egg’ problem: an agent cannot determine if its features are redundant without first knowing what other agents intend to transmit.
To address these gaps, we propose JigsawComm, an end-to-end semantic-aware CP framework that learns to “assemble the puzzle” of multi-agent feature transmission. Our approach is based on the joint semantic feature encoding and transmission problem that we formulate. First, a regularized feature encoder is trained to extract features that are both semantically important and sparse. Second, a lightweight Feature Utility Estimator (FUE) head is introduced to predict the data utility of each agent’s feature to the final perception result, producing compact meta utility maps that (i) approximate the marginal contribution of each agent’s features in the BEV grid to the downstream task and (ii) expose cross-agent feature overlap. This entire framework is trained end-to-end, using a differentiable scheduler to jointly optimize the feature encoder and the FUE. During inference, agents exchange the meta utility maps and locally compute an explainable, deterministic transmission policy: select at most one agent with the highest utility per cell, and greedily admit cells until the communication budget is met, which is provably optimal under our optimization formulation.
JigsawComm’s transmission policy inherently eliminates cross-agent duplication by construction with its top-$1$ per-cell rule. This makes the system practical and scalable: because the final transmission is a sparse ‘jigsaw’ assembled from the best cell-level features across the network, the total communication cost is bounded by the size of a single sparse feature map. This achieves an asymptotic transmission cost of $\mathcal{O}(1)$ with respect to the number of agents $N$, rather than the $\mathcal{O}(N)$ of naive broadcast. The scheduler requires only the exchanged meta utility maps and yields identical decisions across all agents, making it deployable even in the challenging decentralized mode of V2X.
Our contributions can be summarized as follows:
  • We propose JigsawComm, a novel communication-efficient CP framework that maximizes the contribution of each transmitted bit to the final perception task, using a jointly trained semantic sparse feature encoder and end-to-end differentiable scheduler.
  • We introduce a learned utility proxy and a deterministic, provably optimal top-1 scheduling policy that, effectively eliminates cross-agent redundancy, achieving scalable $\mathcal{O}(1)$ communication cost.
  • We extensively evaluate JigsawComm on the OPV2V and DAIR-V2X benchmarks, demonstrating comparable or superior accuracy while reducing communication data volume and improving utility-per-byte both by up to over $500\times$ compared to prior methods [xu2022opv2v, f_cooper, xu2022v2x, hu2022where2comm], setting the new SOTA for communication-efficient CP. [^1] Code available at https://github.com/WiSeR-Lab/JigsawComm.
  • Our results uncover a counterintuitive ‘more is worse’ effect in existing CP fusion backbones, showing that redundant multi-agent features can degrade accuracy rather than improve it. We highlight the need for future CP fusion architectures that can effectively leverage, instead of suffer from, redundancy.
2 Background and Related Work
2.1 Cooperative Perception Paradigms
Research in CP has explored various strategies for fusing information from multiple agents, which are primarily categorized by the stage at which fusion occurs. Early fusion involves the transmission and aggregation of raw sensor data, such as LiDAR point clouds or camera images [zhang2021emp, zhang2023robust]. While this approach theoretically preserves the maximum amount of information, it is prohibitively expensive in terms of required communication bandwidth and is thus considered impractical for real-world V2X applications [f_cooper]. At the other extreme, late fusion operates on the final perception outputs, where agents directly exchange local perception results [rauch2012car2x]. This method is highly communication-efficient but suffers from significant information loss during local processing, making it difficult to resolve detection conflicts and susceptible to error propagation [xu2022opv2v].
The dominant paradigm in modern CP is intermediate fusion, which strikes a balance between these two extremes [f_cooper, xu2022opv2v, hu2022where2comm, xu2022v2x, liu2020when2com, xu2025cosdh]. In this approach, agents first process their raw sensor data into an intermediate feature representation, typically within a shared and aligned BEV grid. These features are then exchanged and fused before being passed to a final detection head. This strategy preserves rich semantic and spatial information while being more communication-friendly than early fusion, forming the basis for most state-of-the-art methods, including JigsawComm proposed in this paper.
2.2 Communication-Efficient CP Strategies
Within the intermediate fusion paradigm, several previous works have focused on optimizing the trade-off between perception performance and communication cost. These efforts can be broadly classified by their core strategy.
Message Representation: Changing How Information is Encoded. This category of methods focuses on compressing the feature maps. While standard encoder-decoder models are common [xu2022opv2v, xu2025cosdh], more advanced strategies have emerged. CodeFilling [hu2024communication, zhao2025quantv2x], for example, uses a shared codebook (i.e. lookup table) to quantize high-dimensional features into compact integer indices. A key limitation of this approach is that the codebook is learned offline and remains static during inference. This makes them difficult to generalize to novel environments. More recently, DiffCP proposed a generative approach [mao2025diffcp], where the sender transmits a small semantic seed, and the receiver leverages a diffusion model to reconstruct the sender’s full BEV map. While achieving very low bitrates, this method incurs a substantial computational burden for the iterative denoising process, violating real-time latency constraints [ADConstraints], and carries an inherent risk of “hallucination”, sacrificing accuracy.
Message Selection: Deciding What to Transmit. This line of research aims to prune the information sent by agents, assuming that not all data is equally valuable. Early work like When2com [liu2020when2com] employed a graph neural network with a handshake mechanism to select which agents to communicate with, performing a coarse, agent-level selection. However, as V2X communication primarily relies on wireless broadcast rather than direct pairwise communications [FCC_DSRC, CV2X_ETSI_EN_303_613], optimizing communication graphs does not save bandwidth in practice. A more fine-grained approach was introduced by Where2comm [hu2022where2comm], which has become the de facto canonical baseline for communication-efficient CP. The idea is to generate a spatial confidence map by applying the detector head on each agent’s local feature. This map serves as a heuristic to identify ‘perceptually critical’ regions for transmission, which achieves region-level sparsity, reducing bandwidth compared to sending dense feature maps. More recent work followed the same idea (e.g[xu2025cosdh] further considered intermediate-late hybrid fusion). However, these methods still rely on heuristic utility proxies, which may not be optimal for the final CP task. They also lack explicit mechanisms to suppress redundant information from multiple collaborators with overlapping observations, resulting in a non-scalable, linear increase of communication cost with respect to the number of agents.
Nevertheless, quantifying and minimizing inter-agent redundancy is a challenging problem. Prior CP schemes that consider redundancy reduction suffer from various limitations. For example, a few early fusion based CP methods adopt geometry-based heuristics (e.g. each agent responsible for partitioned areas that are closest to itself) [zhang2021emp, autocast, wang2024edge, yu2023flow], which ignore sensory data content and semantics (e.g. occlusion), leading to potentially inaccurate results. Also, intermediate fusion-based CP methods that adopt mutual information-based metrics for redundancy reduction are computationally impractical with limited samples [su2024makes].
Table 1: Comparison with existing methods.
[…]
    Refers to whether the method can meet the bandwidth constraint of V2X and is scalable.
  • $\dagger$
    Refers to a method’s ability to adapt to variations in available bandwidth.
Positioning Our Work. The existing landscape of communication-efficient CP reveals a clear set of unresolved challenges: a reliance on heuristic and non-task-optimal proxies for task utility, a lack of explicit modeling and removal of semantic inter-agent redundancy, and static performance-bandwidth trade-offs imposed by fixed representation schemes. Our work addresses these gaps by introducing a semantic-aware, task-oriented, and end-to-end learned scheduling policy that is generalizable, adaptive, and efficiently manages shared communication resources among multiple agents. Table 1 provides a comparative summary.
Figure 2: Overview of JigsawComm.
3 Methodology: End-to-End Communication-Aware Cooperative Perception
3.1 Problem Formulation
We consider $N$ agents $\mathcal{A}=\{a_{1},\dots,a_{N}\}$ (CAVs or RSUs, within each other’s communication range) collaborating via V2X. Per cycle, agent $a_{i}$ obtains raw sensor data $X_{i}$ and uses an encoder $\Phi_{\text{enc}}$ to produce a BEV feature map $\mathbf{F}_{i}=\Phi_{\text{enc}}(X_{i})\in\mathbb{R}^{C\times H\times W}$ with cells $\mathbf{f}_{i}^{l}\in\mathbb{R}^{C}$, $l\in\mathcal{L}=\{1,\dots,L\}$, L=HWL{=}HW. After exchanging meta utility maps ${\mathbf{U}}_{i=1}^{N}$, each agent $a_{i}$ then transmits a sparse subset $\mathbf{F}^{\prime}_{i}=\mathbf{M}_{i}\odot\mathbf{F}_{i}$, where $\mathbf{M}_{i}\in\{0,1\}^{H\times W}$ and $\odot$ denotes element-wise multiplication. A fusion module $\Phi_{\text{fuse}}$ and decoder $\Phi_{\text{dec}}$ produce the output $\hat{Y}$. Our goal is to maximize the CP accuracy within a total communication budget BB [^2] In practice, $B$ equals the available bandwidth $BW$ divided by the frame rate $r$ (e.g. if $BW=20$ Mbps and $r=10$ FPS, $B=2$ Mb).:
$\max_{\theta,\{\mathbf{M}_{i}\}_{i=1}^{N}}$ $\mathcal{P}\big(\Phi_{\text{dec}}(\Phi_{\text{fuse}}(\{\mathbf{M}_{i}\odot\Phi_{\text{enc}}(X_{i})\}_{i=1}^{N})),Y\big)$ (1)
s.t. $\sum_{i=1}^{N}|\mathbf{M}_{i}\odot\Phi_{\text{enc}}(X_{i})|\leq B.$ (2)
Here, $\mathcal{P}(\cdot)$ is the CP accuracy evaluation metric (e.g. mAP), $Y$ corresponds to the ground truth supervision, $\theta$ denotes the trainable parameters for the entire network, and the binary masks $\{\mathbf{M}_{i}\}_{i=1}^{N}$ select subsets of the encoded features across spatial dimensions for transmission and fusion.
Challenges. This formulation presents three core challenges: (C1) During inference, since there is no ground truth, the CP accuracy $\mathcal{P}$ cannot be directly evaluated by the agents. Thus, we must design a utility proxy that is aligned with the CP accuracy and can be computed locally. Previous work uses heuristic local confidence scores [liu2020when2com, hu2022where2comm, sheng2024semantic, xu2025cosdh], which fails to capture cross-agent redundancy and is not necessarily the optimal representations of the final accuracy. (C2) Ideally, a utility proxy needs to be jointly trained with the network under the accuracy objective to guarantee a direct alignment. However, the binary scheduling masks are non-differentiable, which hinders end-to-end training. (C3) Even if the utility function can be computed or approximated, the feature selection process under Eq. (2) is combinatorial, as it reduces to a 0-1 knapsack problem, which is NP-Hard [kellerer2004multidimensional]. Moreover, ‘more bits’ is not better: redundant views congest the V2X data channel without boosting fused utility. Thus, we must transmit only semantically meaningful, non-redundant information and maximize the contribution of every transmitted bit to CP accuracy.
System Overview. JigsawComm learns to assemble the ‘jigsaw’ of multi-agent BEV features so that the CP messages carry only essential, non-overlapping content. As shown in Fig. 2, each agent: 1 produces sparse and semantically important features (Sec. 3.2); 2 predicts a compact meta utility map using a Feature Utility Estimator (FUE) network, which is exchanged among agents efficiently (Sec. 3.3); 3 locally applies a deterministic, redundancy-aware top-1-per-cell policy, producing a feature selection mask that maximizes the total utility while satisfying the communication budget (Sec. 3.4); 4 broadcasts the selected features; 5 fuses the exchanged features and passes to the inference head for output. With the total per-frame communication cost upper-bounded by the size of a single sparse semantic feature, the required bandwidth is $\mathcal{O}(1)$ as the number of agents increases, ensuring high scalability.
3.2 Sparse Semantic Feature Encoder
To ensure that features carry compact and essential semantics using a minimal number of bits, we first shape the encoder’s output to reduce intra-agent redundancy. We apply a learnable threshold $\kappa$ and a semantic loss consisting of an $L_{1}$ regularization on the encoded feature:
$\mathcal{L}_{\text{semantic}}=\frac{1}{NL}\sum_{i=1}^{N}\sum_{l=1}^{L}\big\|\mathbf{f}_{i}^{l}\odot\mathbb{I}\{\mathbf{f}_{i}^{l}>\kappa\}\big\|_{1}.$ (3)
The $L_{1}$ regularization, when applied to model parameters, is known to produce sparse coefficients, suppress irrelevant features, and shape the coefficient amplitudes based on their quantitative contribution to the final outcome [ng2004feature]. However, applying $L_{1}$ directly to raw activations $\|\mathbf{f}_{i}^{l}\|_{1}$ only encourages small magnitudes, yet under stochastic normalization and noise, almost never yields exact zeros; gradients remain nonzero almost everywhere, so uninformative activations linger at small but nonzero values—still consuming bits. Therefore, we incorporate the learnable threshold $\kappa$, which explicitly introduces sparsity.
3.3 Feature Utility Estimator (FUE)
Utility Proxy. To address C1, we propose to maximize a utility function $U(\cdot)$. It represents the total contribution of all agents’ transmitted data, expressed as the sum of individual agent’s feature utility minus their pairwise redundancy. Essentially, our utility function can be regarded as a proxy for the overall information about the CP inference result (on the fused feature) that is contained within the transmitted data. The greater this mutual information is, the higher the perception accuracy [qin2021semantic, strinati2021_6g].
$U^{l}=\underbrace{\sum_{i=1}^{N}m_{i}^{l}\,u_{i}^{l}}_{\text{importance}}-\underbrace{\sum_{j\neq i}\frac{m_{i}^{l}m_{j}^{l}\,\tilde{u}_{ij}^{l}}{2}}_{\text{overlap penalty}},\\ \tilde{u}_{ij}^{l}\triangleq\min\{u_{i}^{l},u_{j}^{l}\}.$ (4)
Here, $m_{i}^{l}\in\{0,1\}$ is the binary transmission decision for $\mathbf{f}_{i}^{l}$, $u_{i}^{l}$ is its estimated importance, representing the amount of information contained in $i$’s feature at location $l$ about the CP inference result. Also, $\tilde{u}_{ij}^{l}$ approximates the pairwise redundancy, upper-bounding the shared information that $i$ and $j$’s features both contain about the CP inference result, where the $\frac{1}{2}$ factor is to avoid double-counting. Finally, the frame-level utility is $U=\sum_{l=1}^{L}U^{l}$.
FUE Head. To estimate the feature utility, we implement FUE as a pointwise linear decoder, similar to the architecture of a classification or regression head $u_{i}^{l}\;=\;\mathrm{ReLU}\!\big(\mathbf{w}^{\top}\mathbf{f}_{i}^{l}+b\big)$, resulting in a sparse meta utility map $\mathbf{U}_{i}\in\mathbb{R}^{1\times H\times W}$. The layer is vectorized as a $1{\times}1$ convolution for efficiency and interpretability, since each location already contains neighboring cells’ information from the convolutional layers in the encoder. The FUE is end-to-end trained with the transmission policy described in Sec. 3.4 and aligned with the perception accuracy loss. Therefore, its value directly relates to the transmission decision such that higher utility values correlate with higher transmission priority and larger contribution to the accuracy.
3.4 Transmission Scheduler
Even after transforming the optimization into maximizing the learned utility proxy, the challenges of non-differentiability in scheduling (C2) and the NP-hard combinatorial knapsack problem (C3) still exist. To solve these, we design a scheduling algorithm that is (i) differentiable at training and (ii) deterministic at inference with provably optimal guarantee given fixed transmission cost per feature.
After exchanging meta utility maps $\{\mathbf{U}_{i}\}_{i=1}^{N}$, the scheduler produces a feature selection mask $\mathbf{M}_{i}$ for each agent based on a straightforward and interpretable policy:
  1. 1.
    For each spatial cell $l$, select at most one agent with the highest non-zero utility $\max_{j}u_{j}^{l}$ to transmit.
  2. 2.
    From this set of candidates, greedily admit cells with the highest utility-to-cost ratio until the budget $B$ is met.
This policy can be deterministically computed by all agents, making it practically compatible with both centralized (V2I) and decentralized (V2V) operation (Sec. 5).
Figure 3: Visualization of meta utility map examples. The top-1 aggregated utility map eliminates redundancy by selecting the agent with the highest utility at each location.
Train-Time Differentiable Scheduling. To address C2, we relax the discrete selection during training. Instead, we model the mask $m_{i}^{l}$ as the composition of two gates:
(1) Importance gate (scalar, per agent–cell). We use a logistic gate with temperature $\eta>0$: $\alpha_{i}^{l}\;=\;\sigma\!\Big(\frac{u_{i}^{l}-\tau}{\eta}\Big)$, where $\sigma(\cdot)$ is the sigmoid function and $\tau$ is a learnable threshold. We anneal $\eta\downarrow 0$ over training to make $\alpha_{i}^{l}$ near-binary.
(2) Cross-agent top-1 (categorical, per cell). To model the “pick one agent per cell” constraint differentiably, we apply the Gumbel–Softmax reparameterization [jang2016categorical] on the vector of scores $\mathbf{u}^{l}=(u_{1}^{l},\dots,u_{N}^{l})$ with an independent Gumbel noise per agent: $g_{i}^{l}\sim\mathrm{Gumbel}(0,1)$ with $g_{i}^{l}=-\log(-\log\varepsilon_{i}^{l})$, $\varepsilon_{i}^{l}\sim\mathrm{Uniform}(0,1)$ , and a temperature $\gamma>0$, resulting in:
$\boldsymbol{\beta}_{i}^{l}(\gamma)=\mathrm{softmax}_{i}\left(\frac{u_{i}^{l}+g_{i}^{l}}{\gamma}\right)=\frac{\exp\left((u_{i}^{l}+g_{i}^{l})/\gamma\right)}{\sum_{k=1}^{N}\exp\left((u_{k}^{l}+g_{k}^{l})/\gamma\right)}.$ (5)
We combine these gates using a Straight-Through Estimator (STE) [liu24differentiable_comb]. In the forward pass, we apply the deterministic policy for stability:
$m_{i}^{l}\_\text{fwd}=\mathbb{I}[u_{i}^{l}\geq\tau]\cdot\text{OneHot}\Big(i=\arg\max_{k}u_{k}^{l}\Big).$ (6)
In the backward pass, we compute gradients as if the mask was the soft, differentiable approximation $\tilde{m}_{i}^{l}=\alpha_{i}^{l}\boldsymbol{\beta}_{i}^{l}(\gamma)$. $\gamma$ follows the same annealing schedule as $\eta$ during training.
Inference-time Deterministic Policy. At inference, we apply the deterministic policy from the STE’s forward pass. Each agent $a_{i}$ applies the thresholding $\tau$ on the meta utility map $\mathbf{U}_{i}$ and exchanges the sparse meta utility map, on which the local scheduler applies the per-cell top-1 policy:
$m_{i}^{l}=\mathbb{I}\!\big[u_{i}^{l}=\max_{k}u_{k}^{l}\big]\cdot\mathbb{I}[u_{i}^{l}\geq\tau].$ (7)
Let $C_{i}^{l}$ be the cost (bytes) of sending $\mathbf{f}_{i}^{l}$, and $\rho_{i}^{l}=u_{i}^{l}/C_{i}^{l}$ the utility–cost ratio. We sort all $(i,l)$ with $m_{i}^{l}=1$ by $\rho_{i}^{l}$ and admit the longest prefix whose cumulative cost $\leq B$.
Proposition 1 (Consistency of differentiable and deterministic schedulers).
For any $(i,l)$ with $u_{i}^{l}\neq\tau$ and with no injected noise at inference, we have
$\lim_{\eta\to 0^{+}}\sigma\!\Big(\frac{u_{i}^{l}-\tau}{\eta}\Big)=\mathbb{I}[u_{i}^{l}>\tau],$ (8)
$\lim_{\gamma\to 0^{+}}\boldsymbol{\beta}^{l}(\gamma)=\mathrm{OneHot}\!\Big(i=\arg\max_{k}u_{k}^{l}\Big).$ (9)
Therefore, we have
$\lim_{\eta\to 0^{+}}\tilde{m}_{i}^{l}=\mathbb{I}[u_{i}^{l}=\max_{k}u_{k}^{l}]\cdot\mathbb{I}[u_{i}^{l}>\tau]=m_{i}^{l}.$ (10)
Prop. (1) indicates that the train-time relaxed mask converges pointwise to the deterministic policy. Also, instead of a heuristic design, we show that this policy is provably optimal for the utility proxy in Eq. (4) and the knapsack problem under uniform costs.
Theorem 1 (Singleton optimality).
For each cell $l$, under the utility proxy in Eq. (4), any maximizer of $U^{l}$ selects at most one agent: $|\{i:m_{i}^{l}=1\}|\leq 1$.
Proof.
Order agents by $u_{1}^{l}\geq u_{2}^{l}\geq\cdots$. Selecting agent $1$ yields $U^{l}=u_{1}^{l}$. Adding agent $2$ changes utility by $u_{2}^{l}-\min(u_{1}^{l},u_{2}^{l})=0$. Adding any $k\geq 3$ to a set $\mathcal{S}$ that already contains agent $1$ and $2$ yields marginal gain $u_{k}^{l}-\sum_{j\in\mathcal{S}}\min(u_{k}^{l},u_{j}^{l})\leq u_{k}^{l}-\min(u_{k}^{l},u_{2}^{l})=0.$
Figure 4: Visualization of feature map examples.
Theorem (1) ensures that there is at most one agent with the highest utility to transmit per location cell, inherently eliminating cross-agent redundancy, as visualized in Fig. 3. This reduces the knapsack problem (C3) of selecting features across both agent and location dimensions to selecting the best set of location cells to transmit. Although the problem remains a knapsack problem, it becomes a uniform-cost knapsack, since the number of channels is fixed per feature, resulting in the same transmission cost. It is known that the uniform-cost knapsack reduces to a cardinality-constrained selection, which can be optimally solved by taking the elements with the largest utilities [kellerer2004multidimensional], as shown in Theorem (2).
Theorem 2 (Greedy optimality under equal costs).
Suppose each feature $(i,l)$ has the same nonzero cost $C_{i}^{l}\equiv c>0$. Let the per-frame budget be $B$ and define the integer capacity $K\triangleq\big\lfloor B/c\big\rfloor$. Then an optimal solution is obtained by selecting any $K$ features with the largest utilities $u_{i}^{l}$.
Proof.
By Theorem (1), each cell yields at most one admissible candidate; thus the global problem is
$\max_{\mathcal{S}\subseteq\mathcal{C}}\ \sum_{(i,l)\in\mathcal{S}}u_{i}^{l}\quad\text{s.t. }\ |\mathcal{S}|\leq K,$ (11)
where $\mathcal{C}$ is the set of at-most-one-per-cell candidates and all costs equal $c$. This is a cardinality-constrained selection problem. Let $\mathcal{G}$ be the greedy solution: the $K$ items with the largest utilities. Assume for contradiction that $\mathcal{G}$ is not optimal. Let $\mathcal{O}$ be an optimal solution with $|\mathcal{O}|\leq K$ and $\sum_{o\in\mathcal{O}}u(o)>\sum_{g\in\mathcal{G}}u(g)$. If $|\mathcal{O}|<K$, augment $\mathcal{O}$ by adding the highest-utility items not in $\mathcal{O}$ until its size reaches $K$; this cannot decrease its objective value because utilities are nonnegative, so w.l.o.g. take $|\mathcal{O}|=K$.
Order items in $\mathcal{G}$ and $\mathcal{O}$ so that $u(g_{1})\geq\cdots\geq u(g_{K})$ and $u(o_{1})\geq\cdots\geq u(o_{K})$. Since $\mathcal{G}$ contains the top-$K$ utilities overall, we must have $u(g_{j})\geq u(o_{j})$ for every $j=1,\dots,K$. Summing over $j$ yields $\sum_{j=1}^{K}u(g_{j})\geq\sum_{j=1}^{K}u(o_{j})$, contradicting $\sum_{o\in\mathcal{O}}u(o)>\sum_{g\in\mathcal{G}}u(g)$. Hence $\mathcal{G}$ is optimal. ∎
3.5 End-to-End Communication-Aware Training
We unify these components into an end-to-end trainable framework. The final loss combines the downstream task loss (e.g. focal, cross-entropy, etc.) with the semantic loss:
$\mathcal{L}_{\text{JigsawComm}}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{semantic}},$ (12)
where $\lambda$ is a hyperparameter balancing the sparsity of the encoded feature and task accuracy. The $\mathcal{L}_{\text{task}}$ gradient flows through the differentiable scheduler (Sec. 3.4) to train the FUE head (Sec. 3.3), while $\mathcal{L}_{\text{semantic}}$ shapes the encoder (Sec. 3.2). By learning the sparsity threshold $\kappa$ and the utility threshold $\tau$ jointly, JigsawComm learns to activate and transmit features only when they contribute to CP accuracy, maximizing the utility of every transmitted bit, as shown in Fig. 4.
4 Evaluation
[…]
Dataset OPV2V DAIR-V2X
Model mAP @0.5 mAP @0.7 Per-Agent Size (KB) Total Data Size (KB) Rel. Eff. (%) mAP @0.5 mAP @0.7 Per-Agent Size (KB) Total Data Size (KB) Rel. Eff. (%)
No Fusion 0.68 0.60 - - - 0.50 0.44 - - -
F-Cooper 0.89 0.79 8,800.00 22,792.00 0.19 0.56 0.39 16,384.00 32,768.00 0.70
V2X-ViT 0.87 0.67 8,800.00 22,792.00 0.18 0.68 0.50 16,384.00 32,768.00 0.84
AttFusion()(*) 0.91 0.82 8,800.00 22,792.00 0.19 0.65 0.51 16,384.00 32,768.00 0.81
()(*) + Enc-Dec 0.89 0.81 4,400.00 11396.00 0.38 0.65 0.51 8,192.00 16,384.00 1.61
Where2comm 0.85 0.60 1,152.47 2,794.72 1.47 0.46 0.23 14,692.88 29,385.76 0.64
JigsawComm 0.92 0.82 17.46 43.84 100.00 0.69 0.52 148.96 281.97 100.00
4.1 Experimental Setup
Dataset. We evaluate JigsawComm on two CP datasets: (i) OPV2V [xu2022opv2v], which is a large-scale benchmark collected across 70 diverse scenes from 8 towns in the digital twin simulator CARLA [Dosovitskiy17Carla]. It includes 11,464 frames of LiDAR point cloud data from 2-5 agents and 232,913 annotated 3D bounding boxes. (ii) DAIR-V2X [yu2022dair], which is the first large-scale real-world CP dataset for vehicle-infrastructure cooperation, consisted of 71,254 frames.
Implementation Details. We adopt PointPillars [lang2019pointpillars] as the LiDAR BEV backbone and a max-out module as the fusion backbone [f_cooper]. We train JigsawComm for up to 100 epochs using Adam ($\beta_{1}\!=\!0.9,\beta_{2}\!=\!0.999$), with a base learning rate of 0.002. The Gumbel-Softmax temperatures $\eta,\gamma$ are annealed from $0.9\to 0.1$. We select $\lambda=1000$ for OPV2V and $\lambda=100$ for DAIR-V2X, through standard cross validation.
Evaluation Metrics. We evaluate 3D object detection performance using the mean Average Precision (mAP) at IoU thresholds of 0.5 and 0.7. Communication cost is measured by the Average Total Data Size (KB) transmitted per frame by all agents, under FP8. We also report Relative Communication Efficiency, defined as $\text{mAP}@0.5/\text{Total Size}$, normalized to our method (100%) for comparison. The required bandwidth is calculated as the total transmitted feature size multiplied by the frame rate ($r=10$ FPS).
Baselines. We follow the evaluation protocol that has become standard in communication-efficient CP [xu2025cosdh, hu2024communication, mao2025diffcp, zhao2025quantv2x]: (i) a no-fusion single-vehicle detector, (ii) strong full-transmission intermediate-fusion methods that transmit dense BEV features, including F-Cooper [f_cooper], V2X-ViT [xu2022v2x], and AttFusion [xu2022opv2v], and (iii) communication-efficient CP via Where2comm [hu2022where2comm], which is the de facto canonical communication-efficient baseline adopted by recent works [xu2025cosdh, hu2024communication, mao2025diffcp, zhao2025quantv2x]. Although some recent approaches [hu2024communication, mao2025diffcp, zhao2025quantv2x] achieve comparable bandwidth requirements, they suffer from practical limitations in generalizability or computation latency, as discussed in Sec. 2.2.
4.2 Quantitative Results
[…]
(a) OPV2V
(b) DAIR-V2X
Figure 5: Tradeoff between accuracy and required bandwidth.
Accuracy-Bandwidth Tradeoff. Figure 5 visualizes JigsawComm’s performance versus the required communication bandwidth ($BW$). The baselines are fixed points, representing inflexible, inefficient strategies. JigsawComm’s curve is generated by varying the communication budget $B$. The accuracy of JigsawComm saturates quickly after $BW\geq$1,000 Kbps, making it a flexible, generalizable, and pragmatic solution for deployment under the limited bandwidth offered by V2X (e.g, $\leq$20 Mbps) [FCC_DSRC, CV2X_ETSI_EN_303_613].
Figure 6: Accuracy, feature size, and required bandwidth grouped by agent counts on OPV2V. Instead of a linear increase in total feature size with more agents, JigsawComm induces an asymptotic $\mathcal{O}(1)$ communication cost while achieving high accuracy.
Scalability with Number of Agents. JigsawComm achieves an asymptotic $\mathcal{O}(1)$ communication cost, where the total bandwidth is bounded by the scene’s information content in a single sparse feature, not the number of participating agents $N$. We validate this claim empirically in Fig. 6 using the OPV2V dataset, which includes 2 to 5 agents. JigsawComm’s accuracy remains high as the number of agents varies. The total feature size grows slightly from $\approx$40 KB at $N=2$ to $\approx$58 KB at $N=4$, which is attributed to the fact that more agents increases collective scene coverage. Our FUE correctly identifies these non-redundant cells as having high utility and adds them to the ‘jigsaw puzzle’.
Figure 7: JigsawComm’s total feature size and required bandwidth under 10 FPS on our custom dataset.
Since OPV2V is limited to 5 agents, we conduct further experiments on a custom-simulated dataset to analyze the asymptotic trend. We utilized the CARLA digital twin simulator [Dosovitskiy17Carla] in conjunction with the OpenCDA framework [opencda] for coordinated multi-agent simulation and data recording. We set up the scene within CARLA’s ‘Town10HD’ map with 10–20 CAVs following realistic traffic rules, at a large multi-lane intersection. The data was collected following the same format and conventions as the public OPV2V benchmark dataset, including sensor configurations, coordinate systems, and CAV parameters, to ensure compatibility with existing CP models and evaluation pipelines. Fig. 7 presents the result of JigsawComm on the custom dataset. We can observe that JigsawComm’s total communication cost does not double when scaling from 10 to 20 agents. It saturates and remains relatively flat when the number of agents is greater than 15, empirically verifying the asymptotic $\mathcal{O}(1)$ communication cost claim. As more agents join, the collective coverage saturates, the overlap becomes dominant, and the marginal contribution of each new agent becomes smaller.
End-to-End Efficiency. Beyond theoretical bandwidth savings, we validate JigsawComm’s end-to-end efficiency. The average end-to-end latency is $45.82\pm 26.96$ ms, which is well within the stringent 100 ms real-time budget for autonomous driving perception [ADConstraints]. This total latency consists of $28.27\pm 23.88$ ms for computation on an RTX 4090 GPU and $17.55\pm 7.14$ ms for communication (assuming a 20 Mbps bandwidth constraint), including the total feature payload from all agents and the meta utility maps with a small total overhead of $0.08\pm 0.04$KB.
4.3 Ablation Study
Table 3: Ablation study of our core components.
Components OPV2V DAIR-V2X
Semantic
Loss
FUE &
Sched.
mAP@0.5 Size (KB) mAP@0.5 Size (KB)
- 0.93 26,292.36 0.55 22,670.62
- 0.94 13,516.80 0.61 7,374.995
0.92 43.84 0.69 279.97
Impact of Core Components. Table 3 ablates our two main contributions: the Semantic Loss $\mathcal{L}_{\text{semantic}}$ and the FUE & Scheduling module. Removing the FUE & Scheduler still allows the encoder’s $\mathcal{L}_{\text{semantic}}$ to create sparse feature maps. However, the entire maps are transmitted, leading to redundancy and a large payload. Without $\mathcal{L}_{\text{semantic}}$, the FUE & Scheduler deal with dense and semantically-overloaded features, making it hard to select due to noisy utility estimates, still resulting in a large payload. Only with the presence of both components, the transmission size is greatly reduced while attaining high CP accuracy.
Effect of Redundancy. The system’s transmission model is based on Theorem (1), demonstrating that selecting features from more than the Top-1 agent per cell is suboptimal. Fig. 8 validates this empirically. Comparing our Top-1 Agent policy to a Top-2 Agents policy, results show that Top-2 is worse at nearly every bandwidth point, indicating that transmitting redundant features wastes bandwidth and degrades perception accuracy. Although adding more data seems beneficial, this highlights a limitation in current fusion networks. Sec. 5 discusses this phenomenon and its implications further.
5 Discussion
On the ‘Cost’ of Redundancy. A critical, counterintuitive finding from our evaluation is that transmitting from the Top-2 agents per cell yields worse performance than transmitting the Top-1 agent alone, as shown in Fig. 8. This ‘more is worse’ outcome stems from a fundamental limitation presented in current SOTA CP fusion algorithms, such as element-wise max-pooling or dot-product attention [xu2022opv2v, hu2022where2comm, f_cooper]. The fused feature is a weighted sum of the received features, with the informative activations in the fused feature upper-bounded by the maximum signal across the individual features. Therefore, unlike human vision, these methods struggle to synergetically fuse redundant information. If the second-best agent’s feature is noisier or semantically different (e.g. conflicting channel activations), it can ‘dilute’ the clean, high-utility representation from the Top-1 agent. This finding corroborates the design philosophy of JigsawComm, which selects only the highest-quality, non-redundant data, resulting in smaller overhead and higher accuracy.
(a) OPV2V
(b) DAIR-V2X
Figure 8: Ablation on number of agents sharing per cell. Including the agent with the second-highest utility does not improve the utility-bandwidth tradeoff, which corroborates Theorem (1).
Why Not a Naive Communication Loss? A straightforward approach to reduce data size might be to add a simple communication loss directly on the transmission masks $\{\mathbf{M}\}_{i=1}^{N}$, penalizing the total number of transmitted bits. We found this approach to produce suboptimal performance in practice. Such a loss disregards the semantic importance and redundancy of the information transmitted, yet penalizes all communication. Our framework avoids this by training the FUE head to learn a task-aligned utility proxy, ensuring that selection decisions are driven by semantic-importance and redundancy.
Relation to Semantic Communications. JigsawComm is related to semantic/goal-oriented communications [strinati2021_6g, binucci2022adaptive, pezone2022goal], since they share the high-level goal of task-oriented data compression. However, existing work in semantic communications targets different applications [strinati2021_6g], considers only single sensor perception [yuan2023scalable, binucci2022adaptive, pezone2022goal], or neglects sensor correlations [sheng2024semantic]. Applying semantic communications to CP involves unique challenges due to its collaborative and multi-source input nature, where inter-agent redundancy prevails due to overlapping sensory data. Essentially, the CP problem studied in this paper can be reformulated as a joint, multi-source, task-oriented data compression problem, which is difficult to solve without knowing a priori the exact inter-agent data correlation structure. JigsawComm solves this problem via a data-driven approach (with a small overhead of exchanging semantic feature utilities).
Meanwhile, the semantic regularization and redundancy-eliminating transmission scheduling design of JigsawComm can be viewed as a way of implementing the information bottleneck (IB) principle, where unnecessary information is minimized while task-relevant information is maximized, which is essential to achieve goal-oriented communications [strinati2021_6g, pezone2022goal]. Interestingly, the same IB principle has also been shown to be effective in reducing generalization errors in deep learning models [kawaguchi2023does]. Here, our end-to-end trained JigsawComm learns to keep only semantically-relevant information from its sparse encoder and selects non-redundant and high-utility features for transmission, which effectively controls the IB of the CP pipeline. Therefore, in addition to the drastic reduction in required bandwidth for real-time communication, JigsawComm also achieves matching or improved perceptual quality compared to full-feature baselines.
Practical Implementation. Our framework is designed for practical V2X implementation by leveraging its dual-channel architecture [FCC_DSRC, CV2X_ETSI_EN_303_613]. The meta utility maps, though scaling linearly with agent count, can be transmitted over the control channel, incurring negligible overhead (<<0.03 KB per agent under 4-bit quantization). The final, sparse BEV features are then sent efficiently over the data channel, such as via the Collective Perception Message (CPM) defined by ITS [cpm]. Note that, our scheduler outputs the transmission priorities of all agents’ selected features as a byproduct (ordered by the corresponding feature utility values), which can be easily integrated into the low-layer V2X protocols for both scheduling-based (e.g. C-V2X, 5G NR-V2X [CV2X_ETSI_EN_303_613, garcia2021tutorial]) and random-access V2X protocols (e.g. DSRC [FCC_DSRC]). Our scheme is compatible with both centralized (V2I) and decentralized (V2V) operating modes. For V2I, an edge server co-located with the RSU can perform scheduling, fusion, and inference. For V2V, the same set of computations is performed by each agent in a distributed manner.
6 Conclusion and Future Work
We present JigsawComm, a novel joint semantic feature encoding and transmission framework for CP that maximizes every transmitted bit’s contribution to the final task accuracy. By jointly training a semantic sparse encoder and a lightweight Feature Utility Estimator via a differentiable scheduler, JigsawComm learns to extract and select only essential, non-redundant data. This enables a provably optimal top-1 policy at inference that eliminates cross-agent redundancy and achieves a scalable $\mathcal{O}(1)$ communication cost. Our experiments on OPV2V and DAIR-V2X demonstrate SOTA accuracy with up to over $500\times$ bandwidth reduction.
There are several promising directions for future work. First, as JigsawComm adopts the index-based encoding to convert sparsely encoded features to bits, we will further push the limit of JigsawComm’s bandwidth efficiency by integrating traditional lossless source coding methods such as entropy coding [storer1988data]. Second, we will also extend JigsawComm with joint semantic source and channel coding to deal with unreliable channel conditions. Third, we will study the impact of different sensor modalities and feature cell granularities on the accuracy-bandwidth tradeoff, and explore novel fusion networks that can leverage, rather than suffer from, redundancy. Finally, we will study practical V2X-based CP system design, robustness to communication failures or packet losses, and end-to-end goal-oriented CP systems for autonomous driving.

### v2

@@
 With its great potential for facilitating more informed driving decisions and enhancing road safety, CP is a critical technology for 50–90% of the vehicle market that CAVs are anticipated to comprise by 2040 [Chatziioannou2024].
-Figure 1: A sample scenario from OPV2V dataset (left).
-Under the same bandwidth limit, prior work [hu2022where2comm] (mid) does not utilize bandwidth efficiently, leading to missed detection and reduced accuracy.
+Figure 1: A sample scenario from the OPV2V dataset (left).
+Under the same bandwidth limit, prior work [hu2022where2comm] (middle) does not utilize bandwidth efficiently, leading to missed detections.
 JigsawComm (right) maximizes every bit’s contribution, enabling complete and accurate detection.
-Quantitative evaluation and discussion on the ‘cost’ of redundancy can be found in Sec.
-4 and Sec.
-5, respectively.
+See Sec.
+4 for quantitative evaluation and discussion on the “cost” of redundancy.
 Despite this promise, CP faces a formidable challenge in communication efficiency.
@@
 <<100 ms [ADConstraints]).
-As previous research also demonstrated that exchanging local detection results alone is suboptimal [xu2022opv2v], the field has largely converged on the intermediate fusion framework—exchanging and fusing Bird’s-Eye-View (BEV) features—as the most effective paradigm [xu2022opv2v].
+As previous research also demonstrated that exchanging local detection results alone is suboptimal [xu2022opv2v], the field has largely converged on the intermediate fusion framework—exchanging and fusing Bird’s-Eye-View (BEV) features—as the most effective paradigm [xu2022opv2v, hu2022where2comm, xu2022v2x, f_cooper].
 Even so, transmitting dense deep features from multiple agents still exhausts V2X bandwidth.
-The key challenge stems from the high degree of data redundancy present in these shared features, which can be classified into two forms: (i) Intra-agent redundancy, where features waste bits on semantically irrelevant information; and (ii) Inter-agent redundancy, where multiple agents transmit overlapping and unnecessary information.
-As shown in Fig.
-1, under limited bandwidth, such inefficient communication not only leads to reduced detection coverage but can even degrade perception accuracy when lower quality data dilutes the utility of the better ones after fusion.
-Existing work on communication-efficient CP fails to solve this.
-They typically adopt one of two strategies: (i) Compression methods, such as encoder-decoders [xu2022opv2v] or codebooks [hu2024communication, zhao2025quantv2x].
-Such methods either only reduce the data size but still transmit the entire feature map, failing to address redundancy, or overly restrict the representation power to offline-trained dictionaries, undermining generalizability; (ii) Message selection methods [liu2020when2com, hu2022where2comm, hu2024communication, xu2025cosdh], choosing ‘critical’ regions to transmit but relying on simple local heuristics (e.g.
-local detection confidence) that are poor proxies for the final cooperative task.
-Crucially, both approaches ignore cross-agent redundancy.
-As a result, they still require significant bandwidth and are not scalable: their communication cost grows linearly with the number of agents.
+The key challenge stems from the high degree of data redundancy in the shared features, which takes two forms: (i) intra-agent redundancy, where features waste bits on semantically irrelevant information; and (ii) inter-agent redundancy, where multiple agents transmit overlapping data about the same scene regions.
+As illustrated in Fig.
+1, under limited bandwidth such inefficiency not only reduces detection coverage but can even degrade perception accuracy when lower-quality data dilutes higher-quality representations after fusion (Sec.
+4.4).
+Existing communication-efficient CP methods fail to address both forms simultaneously.
+They typically adopt one of two strategies: (i) Compression, using encoder-decoders [xu2022opv2v] or codebooks [hu2024communication, zhao2025quantv2x], which reduce data size but still transmit the entire feature map or restrict representation power to offline-trained dictionaries; (ii) Message selection [liu2020when2com, hu2022where2comm, hu2024communication, xu2025cosdh], choosing “critical” regions based on local heuristics (e.g.
+detection confidence) that loosely correlate with the final cooperative task.
+Crucially, both approaches ignore cross-agent redundancy, so their communication cost still grows linearly with the number of agents.
 The problem of designing a truly communication-efficient, scalable, and accurate CP system remains open.
-To design such a CP system, we argue that it must maximize the contribution of every transmitted bit towards the final perception goal.
-This requires that only essential, semantically meaningful, and non-redundant information is transmitted.
-Therefore, the feature encoding (local compression) and transmission (message selection) algorithms must be jointly designed to realize this potential.
-This joint design is non-trivial.
-Removing semantic, cross-agent redundancy is a challenging ‘chicken-and-egg’ problem: an agent cannot determine if its features are redundant without first knowing what other agents intend to transmit.
+We argue that a practical CP system must maximize the contribution of every transmitted bit to the final perception goal, transmitting only essential, semantically meaningful, and non-redundant information.
+This requires jointly designing the feature encoding and transmission selection, which is non-trivial: removing cross-agent redundancy is a “chicken-and-egg” problem, since an agent cannot determine whether its features are redundant without knowing what other agents intend to transmit.
 To address these gaps, we propose JigsawComm, an end-to-end semantic-aware CP framework that learns to “assemble the puzzle” of multi-agent feature transmission.
-Our approach is based on the joint semantic feature encoding and transmission problem that we formulate.
-First, a regularized feature encoder is trained to extract features that are both semantically important and sparse.
-Second, a lightweight Feature Utility Estimator (FUE) head is introduced to predict the data utility of each agent’s feature to the final perception result, producing compact meta utility maps that (i) approximate the marginal contribution of each agent’s features in the BEV grid to the downstream task and (ii) expose cross-agent feature overlap.
-This entire framework is trained end-to-end, using a differentiable scheduler to jointly optimize the feature encoder and the FUE.
-During inference, agents exchange the meta utility maps and locally compute an explainable, deterministic transmission policy: select at most one agent with the highest utility per cell, and greedily admit cells until the communication budget is met, which is provably optimal under our optimization formulation.
-JigsawComm’s transmission policy inherently eliminates cross-agent duplication by construction with its top-$1$ per-cell rule.
-This makes the system practical and scalable: because the final transmission is a sparse ‘jigsaw’ assembled from the best cell-level features across the network, the total communication cost is bounded by the size of a single sparse feature map.
-This achieves an asymptotic transmission cost of $\mathcal{O}(1)$ with respect to the number of agents $N$, rather than the $\mathcal{O}(N)$ of naive broadcast.
-The scheduler requires only the exchanged meta utility maps and yields identical decisions across all agents, making it deployable even in the challenging decentralized mode of V2X.
-Our contributions can be summarized as follows:
-  • We propose JigsawComm, a novel communication-efficient CP framework that maximizes the contribution of each transmitted bit to the final perception task, using a jointly trained semantic sparse feature encoder and end-to-end differentiable scheduler.
-  • We introduce a learned utility proxy and a deterministic, provably optimal top-1 scheduling policy that, effectively eliminates cross-agent redundancy, achieving scalable $\mathcal{O}(1)$ communication cost.
-  • We extensively evaluate JigsawComm on the OPV2V and DAIR-V2X benchmarks, demonstrating comparable or superior accuracy while reducing communication data volume and improving utility-per-byte both by up to over $500\times$ compared to prior methods [xu2022opv2v, f_cooper, xu2022v2x, hu2022where2comm], setting the new SOTA for communication-efficient CP.
-[^1] Code available at https://github.com/WiSeR-Lab/JigsawComm.
+First, a regularized feature encoder extracts features that are both semantically important and sparse.
+Second, a lightweight Feature Utility Estimator (FUE) predicts each agent’s per-cell contribution to the final perception result, producing compact meta utility maps that approximate the marginal value of each feature and expose cross-agent overlap.
+The entire framework is trained end-to-end via a differentiable scheduler that jointly optimizes the encoder and FUE.
+At inference, agents exchange meta utility maps and locally compute an explainable, deterministic transmission policy: select at most one agent with the highest utility per cell, then greedily admit cells until the communication budget is met.
+We prove this policy is optimal under the learned utility proxy (Theorems 3.1 and 3.2) and validate its alignment with task accuracy empirically.
+The top-1-per-cell rule inherently eliminates cross-agent duplication by construction, making the system practical and scalable: the final transmission is a sparse “jigsaw” assembled from the best cell-level features across agents, bounding the data-channel cost to $\mathcal{O}(1)$ with respect to the number of agents $N$, compared to $\mathcal{O}(N)$ for naive broadcast.
+Although exchanging the lightweight meta utility maps incurs $\mathcal{O}(N)$ signaling, this overhead is negligible in practice (<<0.03 KB per frame).
+The scheduler yields identical decisions on every agent, making it deployable in both centralized (V2I) and decentralized (V2V) modes.
+Our contributions are:
+  • We propose JigsawComm, a CP framework that jointly trains a sparse semantic encoder and an end-to-end differentiable scheduler to maximize each transmitted bit’s contribution to perception accuracy.
+  • We introduce a learned, task-aligned utility proxy and a top-1-per-cell scheduling policy that eliminates cross-agent redundancy and bounds data-channel cost to $\mathcal{O}(1)$ with respect to the number of agents.
+[…]
   • Our results uncover a counterintuitive ‘more is worse’ effect in existing CP fusion backbones, showing that redundant multi-agent features can degrade accuracy rather than improve it.
@@
 These efforts can be broadly classified by their core strategy.
-Message Representation: Changing How Information is Encoded.
-This category of methods focuses on compressing the feature maps.
-While standard encoder-decoder models are common [xu2022opv2v, xu2025cosdh], more advanced strategies have emerged.
-CodeFilling [hu2024communication, zhao2025quantv2x], for example, uses a shared codebook (i.e.
-lookup table) to quantize high-dimensional features into compact integer indices.
-A key limitation of this approach is that the codebook is learned offline and remains static during inference.
-This makes them difficult to generalize to novel environments.
-More recently, DiffCP proposed a generative approach [mao2025diffcp], where the sender transmits a small semantic seed, and the receiver leverages a diffusion model to reconstruct the sender’s full BEV map.
-While achieving very low bitrates, this method incurs a substantial computational burden for the iterative denoising process, violating real-time latency constraints [ADConstraints], and carries an inherent risk of “hallucination”, sacrificing accuracy.
-Message Selection: Deciding What to Transmit.
+Message Representation.
+This category compresses feature maps via encoder-decoders[xu2022opv2v, xu2025cosdh, Zhang_2024_CVPR], codebook quantization [hu2024communication, zhao2025quantv2x], or generative reconstruction [Zhang_2024_CVPR, mao2025diffcp, zhou2025pragmatic].
+Encoder-decoder approaches apply a bottleneck to reduce the feature dimensionality, achieving fixed compression ratios but lacking adaptability to varying scene complexity.
+Codebooks approximate continuous features with discrete indices matching to an offline trained dictionary, which may restrict the generalizability of the representation [hu2024communication, zhao2025quantv2x].
+Diffusion-based methods achieve high accuracy-bandwidth tradeoff (e.g.
+0.86 mAP@0.5 under 80 Kbps [mao2025diffcp]) but the multi-step denoising violate real-time computation latency constraints for autonomous driving [ADConstraints].
+Crucially, these approaches compress the entire feature map uniformly without considering which regions are task-relevant, wasting bits on uninformative background.
+Message Selection.
 This line of research aims to prune the information sent by agents, assuming that not all data is equally valuable.
-Early work like When2com [liu2020when2com] employed a graph neural network with a handshake mechanism to select which agents to communicate with, performing a coarse, agent-level selection.
+Early work like When2com [liu2020when2com] employed a GNN to select which agents to communicate with, performing a coarse, agent-level selection.
 However, as V2X communication primarily relies on wireless broadcast rather than direct pairwise communications [FCC_DSRC, CV2X_ETSI_EN_303_613], optimizing communication graphs does not save bandwidth in practice.
-A more fine-grained approach was introduced by Where2comm [hu2022where2comm], which has become the de facto canonical baseline for communication-efficient CP.
-The idea is to generate a spatial confidence map by applying the detector head on each agent’s local feature.
+A more fine-grained approach was introduced by Where2comm [hu2022where2comm], which has become the de facto canonical baseline for communication-efficient CP, which generate a spatial confidence map by applying the detector head on each agent’s local feature.
 This map serves as a heuristic to identify ‘perceptually critical’ regions for transmission, which achieves region-level sparsity, reducing bandwidth compared to sending dense feature maps.
-More recent work followed the same idea (e.g[xu2025cosdh] further considered intermediate-late hybrid fusion).
-However, these methods still rely on heuristic utility proxies, which may not be optimal for the final CP task.
-They also lack explicit mechanisms to suppress redundant information from multiple collaborators with overlapping observations, resulting in a non-scalable, linear increase of communication cost with respect to the number of agents.
-Nevertheless, quantifying and minimizing inter-agent redundancy is a challenging problem.
-Prior CP schemes that consider redundancy reduction suffer from various limitations.
-For example, a few early fusion based CP methods adopt geometry-based heuristics (e.g.
-each agent responsible for partitioned areas that are closest to itself) [zhang2021emp, autocast, wang2024edge, yu2023flow], which ignore sensory data content and semantics (e.g.
-occlusion), leading to potentially inaccurate results.
-Also, intermediate fusion-based CP methods that adopt mutual information-based metrics for redundancy reduction are computationally impractical with limited samples [su2024makes].
-Table 1: Comparison with existing methods.
-[…]
-    Refers to whether the method can meet the bandwidth constraint of V2X and is scalable.
-  • $\dagger$
-    Refers to a method’s ability to adapt to variations in available bandwidth.
+Follow-up work extends this idea with spatiotemporal fusion [tang2025cost] or intermediate-late hybrid fusion [xu2025cosdh].
+However, these methods rely on local detection confidence that may not correlate well with the final CP accuracy and lack mechanisms to suppress inter-agent redundancy, resulting in $\mathcal{O}(N)$ communication scaling.
+Prior attempts at redundancy reduction either use geometry-based heuristics that ignore semantics [zhang2021emp, autocast, wang2024edge, yu2023flow] or mutual information metrics that are computationally impractical [su2024makes, wei2025infocom].
 Positioning Our Work.
-The existing landscape of communication-efficient CP reveals a clear set of unresolved challenges: a reliance on heuristic and non-task-optimal proxies for task utility, a lack of explicit modeling and removal of semantic inter-agent redundancy, and static performance-bandwidth trade-offs imposed by fixed representation schemes.
-Our work addresses these gaps by introducing a semantic-aware, task-oriented, and end-to-end learned scheduling policy that is generalizable, adaptive, and efficiently manages shared communication resources among multiple agents.
-Table 1 provides a comparative summary.
+Existing communication-efficient CP methods share three unresolved shortcomings: reliance on heuristic utility proxies; no explicit modeling of semantic inter-agent redundancy; and static performance-bandwidth trade-offs.
+Our work addresses all three by introducing a semantic-aware, task-oriented, end-to-end learned scheduling policy that is adaptive to varying bandwidth.
+While JigsawComm shares the high-level goal of task-oriented compression with semantic/goal-oriented communications [strinati2021_6g, pezone2022goal], existing work in that field targets single-source scenarios [yuan2023scalable, binucci2022adaptive] or neglects inter-sensor correlations [sheng2024semantic].
+CP is inherently a multi-source problem where inter-agent redundancy must be modeled without knowing the data correlation structure a priori; JigsawComm solves this in a data-driven manner, with minimal overhead.
 Figure 2: Overview of JigsawComm.
-3 Methodology: End-to-End Communication-Aware Cooperative Perception
+Each agent encodes sparse BEV features, estimates a meta utility map via the FUE, exchanges utility maps, and applies the deterministic top-1-per-cell scheduler to select non-redundant features for transmission and fusion.
+3 Methodology
 3.1 Problem Formulation
-We consider $N$ agents $\mathcal{A}=\{a_{1},\dots,a_{N}\}$ (CAVs or RSUs, within each other’s communication range) collaborating via V2X.
+We consider $N$ agents $\mathcal{A}=\{a_{1},\dots,a_{N}\}$ collaborating via V2X [^2] Here, agents can be CAVs or RSUs, within the same V2X communication range..
 Per cycle, agent $a_{i}$ obtains raw sensor data $X_{i}$ and uses an encoder $\Phi_{\text{enc}}$ to produce a BEV feature map $\mathbf{F}_{i}=\Phi_{\text{enc}}(X_{i})\in\mathbb{R}^{C\times H\times W}$ with cells $\mathbf{f}_{i}^{l}\in\mathbb{R}^{C}$, $l\in\mathcal{L}=\{1,\dots,L\}$, L=HWL{=}HW.
@@
 A fusion module $\Phi_{\text{fuse}}$ and decoder $\Phi_{\text{dec}}$ produce the output $\hat{Y}$.
-Our goal is to maximize the CP accuracy within a total communication budget BB [^2] In practice, $B$ equals the available bandwidth $BW$ divided by the frame rate $r$ (e.g.
-if $BW=20$ Mbps and $r=10$ FPS, $B=2$ Mb).:
+Our goal is to maximize the CP accuracy within a total communication budget BB [^3] BB equals the available bandwidth $BW$ divided by the frame rate $r$.:
 $\max_{\theta,\{\mathbf{M}_{i}\}_{i=1}^{N}}$ $\mathcal{P}\big(\Phi_{\text{dec}}(\Phi_{\text{fuse}}(\{\mathbf{M}_{i}\odot\Phi_{\text{enc}}(X_{i})\}_{i=1}^{N})),Y\big)$ (1)
@@
 mAP), $Y$ corresponds to the ground truth supervision, $\theta$ denotes the trainable parameters for the entire network, and the binary masks $\{\mathbf{M}_{i}\}_{i=1}^{N}$ select subsets of the encoded features across spatial dimensions for transmission and fusion.
+During training, Eq.
+(1) is optimized jointly over the model parameters $\theta$ and the transmission masks $\{\mathbf{M}_{i}\}$ via the differentiable scheduler (Sec.
+3.4); at inference, $\theta$ is fixed and only the scheduling masks $\{\mathbf{M}_{i}\}$ are optimized.
 Challenges.
-This formulation presents three core challenges: (C1) During inference, since there is no ground truth, the CP accuracy $\mathcal{P}$ cannot be directly evaluated by the agents.
-Thus, we must design a utility proxy that is aligned with the CP accuracy and can be computed locally.
-Previous work uses heuristic local confidence scores [liu2020when2com, hu2022where2comm, sheng2024semantic, xu2025cosdh], which fails to capture cross-agent redundancy and is not necessarily the optimal representations of the final accuracy.
-(C2) Ideally, a utility proxy needs to be jointly trained with the network under the accuracy objective to guarantee a direct alignment.
-However, the binary scheduling masks are non-differentiable, which hinders end-to-end training.
-(C3) Even if the utility function can be computed or approximated, the feature selection process under Eq.
-(2) is combinatorial, as it reduces to a 0-1 knapsack problem, which is NP-Hard [kellerer2004multidimensional].
-Moreover, ‘more bits’ is not better: redundant views congest the V2X data channel without boosting fused utility.
-Thus, we must transmit only semantically meaningful, non-redundant information and maximize the contribution of every transmitted bit to CP accuracy.
+This formulation presents three core challenges: (C1) During inference, each agent must decide which features to transmit without access to ground truth or full features from other agents (infeasible to transmit) to optimize for $\mathcal{P}(\cdot)$.
+Thus, a locally computable and compact utility proxy that correlates with CP accuracy and captures cross-agent redundancy is required.
+(C2) The binary scheduling masks $\mathbf{M}_{i}$ are non-differentiable, preventing direct end-to-end training of the utility proxy.
+(C3) The feature selection under Eq.
+(2) is a 0-1 knapsack problem, which is NP-Hard in general [kellerer2004multidimensional].
 System Overview.
@@
 3.4); 4 broadcasts the selected features; 5 fuses the exchanged features and passes to the inference head for output.
-With the total per-frame communication cost upper-bounded by the size of a single sparse semantic feature, the required bandwidth is $\mathcal{O}(1)$ as the number of agents increases, ensuring high scalability.
+With the total per-frame data payload upper-bounded by the size of a single sparse semantic feature, the data-channel bandwidth scales as $\mathcal{O}(1)$ with respect to the number of agents, ensuring high scalability.
 3.2 Sparse Semantic Feature Encoder
 To ensure that features carry compact and essential semantics using a minimal number of bits, we first shape the encoder’s output to reduce intra-agent redundancy.
-We apply a learnable threshold $\kappa$ and a semantic loss consisting of an $L_{1}$ regularization on the encoded feature:
+We apply a semantic loss consisting of an $L_{1}$ regularization and a learnable threshold $\kappa$ on the encoded feature $\mathbf{f}_{i}$:
 $\mathcal{L}_{\text{semantic}}=\frac{1}{NL}\sum_{i=1}^{N}\sum_{l=1}^{L}\big\|\mathbf{f}_{i}^{l}\odot\mathbb{I}\{\mathbf{f}_{i}^{l}>\kappa\}\big\|_{1}.$ (3)
-The $L_{1}$ regularization, when applied to model parameters, is known to produce sparse coefficients, suppress irrelevant features, and shape the coefficient amplitudes based on their quantitative contribution to the final outcome [ng2004feature].
-However, applying $L_{1}$ directly to raw activations $\|\mathbf{f}_{i}^{l}\|_{1}$ only encourages small magnitudes, yet under stochastic normalization and noise, almost never yields exact zeros; gradients remain nonzero almost everywhere, so uninformative activations linger at small but nonzero values—still consuming bits.
-Therefore, we incorporate the learnable threshold $\kappa$, which explicitly introduces sparsity.
+$L_{1}$ regularization is known to produce sparse coefficients and suppress irrelevant features [ng2004feature].
+However, applying it directly to raw activations ($\|\mathbf{f}_{i}^{l}\|_{1}$) only encourages small magnitudes without yielding exact zeros—uninformative activations linger at small but nonzero values, still consuming bits.
+The learnable threshold $\kappa$ resolves this by explicitly zeroing out activations below $\kappa$, producing structurally sparse features with exact zeros that incur no transmission cost under sparse index encoding [hu2022where2comm].
 3.3 Feature Utility Estimator (FUE)
 Utility Proxy.
-To address C1, we propose to maximize a utility function $U(\cdot)$.
-It represents the total contribution of all agents’ transmitted data, expressed as the sum of individual agent’s feature utility minus their pairwise redundancy.
-Essentially, our utility function can be regarded as a proxy for the overall information about the CP inference result (on the fused feature) that is contained within the transmitted data.
-The greater this mutual information is, the higher the perception accuracy [qin2021semantic, strinati2021_6g].
+As established in task-oriented and semantic communication [qin2021semantic, strinati2021_6g], higher CP accuracy correlates with transmitted data retaining more information about the final inference result.
+Therefore, to address (C1), we design a utility function $U(\cdot)$ that serves as a proxy for the downstream CP accuracy and can be computed locally from the agents’ features.
+$U$ aggregates the individual feature contributions of all agents and penalizes pairwise redundancy.
+The utility at each spatial cell $l$ is defined as:
 $U^{l}=\underbrace{\sum_{i=1}^{N}m_{i}^{l}\,u_{i}^{l}}_{\text{importance}}-\underbrace{\sum_{j\neq i}\frac{m_{i}^{l}m_{j}^{l}\,\tilde{u}_{ij}^{l}}{2}}_{\text{overlap penalty}},\\ \tilde{u}_{ij}^{l}\triangleq\min\{u_{i}^{l},u_{j}^{l}\}.$ (4)
-Here, $m_{i}^{l}\in\{0,1\}$ is the binary transmission decision for $\mathbf{f}_{i}^{l}$, $u_{i}^{l}$ is its estimated importance, representing the amount of information contained in $i$’s feature at location $l$ about the CP inference result.
-Also, $\tilde{u}_{ij}^{l}$ approximates the pairwise redundancy, upper-bounding the shared information that $i$ and $j$’s features both contain about the CP inference result, where the $\frac{1}{2}$ factor is to avoid double-counting.
-Finally, the frame-level utility is $U=\sum_{l=1}^{L}U^{l}$.
+Here, $m_{i}^{l}\in\{0,1\}$ is the binary transmission decision for $\mathbf{f}_{i}^{l}$ and $u_{i}^{l}$ is its estimated importance—the FUE predicted contribution of agent $i$’s feature at cell $l$ to the CP result.
+The term $\tilde{u}_{ij}^{l}$ conservatively upper-bounds the pairwise redundancy between agents $i$ and $j$, with the $\frac{1}{2}$ factor avoiding double-counting [cover1999elements].
+Since features in the same BEV cell describe the same physical location; overlapping observations typically differ only in strength, making $\min(\cdot)$ a tractable upper bound on shared information.
+The frame-level utility is $U=\sum_{l=1}^{L}U^{l}$, which serves as a first-order approximation of task-relevant information by rewarding informative features and penalizing redundancy [cover1999elements].
+Scope of the Utility Proxy.
+We emphasize that $U$ is not an exact information-theoretic quantity but a learned, task-aligned intermediate representation acting as an inductive bias that facilitates transmission scheduling: the FUE predicted utilities $u_{i}^{l}$ are trained end-to-end to correlate with CP accuracy in a data-driven fashion rather than derived from a closed-form mutual information expression, which is computationally infeasible [cover1999elements].
+The theoretical guarantees in Sec.
+3.4 (Theorems 3.13.2) hold with respect to this learned proxy; their practical value depends on how well $U$ tracks the true task performance, which we validate empirically in Sec.
+4.
 FUE Head.
-To estimate the feature utility, we implement FUE as a pointwise linear decoder, similar to the architecture of a classification or regression head $u_{i}^{l}\;=\;\mathrm{ReLU}\!\big(\mathbf{w}^{\top}\mathbf{f}_{i}^{l}+b\big)$, resulting in a sparse meta utility map $\mathbf{U}_{i}\in\mathbb{R}^{1\times H\times W}$.
-The layer is vectorized as a $1{\times}1$ convolution for efficiency and interpretability, since each location already contains neighboring cells’ information from the convolutional layers in the encoder.
-The FUE is end-to-end trained with the transmission policy described in Sec.
-3.4 and aligned with the perception accuracy loss.
-Therefore, its value directly relates to the transmission decision such that higher utility values correlate with higher transmission priority and larger contribution to the accuracy.
+We implement the FUE as a single $1{\times}1$ convolution (pointwise linear decoder): $u_{i}^{l}=\mathrm{ReLU}\!\big(\mathbf{w}^{\top}\mathbf{f}_{i}^{l}+b\big)$, producing a sparse meta utility map $\mathbf{U}_{i}\in\mathbb{R}^{1\times H\times W}$.
+Each cell’s receptive field already spans its spatial neighborhood through the encoder’s convolutional layers, so a pointwise head suffices.
+The FUE is trained end-to-end through the differentiable scheduler (Sec.
+3.4): the task loss gradient back-propagates through the scheduling decisions to the FUE weights, aligning higher utility values with higher transmission priority and greater contribution to accuracy.
 3.4 Transmission Scheduler
 Even after transforming the optimization into maximizing the learned utility proxy, the challenges of non-differentiability in scheduling (C2) and the NP-hard combinatorial knapsack problem (C3) still exist.
-To solve these, we design a scheduling algorithm that is (i) differentiable at training and (ii) deterministic at inference with provably optimal guarantee given fixed transmission cost per feature.
+To solve these, we design a scheduling algorithm that is $(i)$ differentiable at training and $(ii)$ deterministic at inference, with a theoretically grounded optimality guarantee under the learned utility proxy for uniform-cost features.
 After exchanging meta utility maps $\{\mathbf{U}_{i}\}_{i=1}^{N}$, the scheduler produces a feature selection mask $\mathbf{M}_{i}$ for each agent based on a straightforward and interpretable policy:
@@
     From this set of candidates, greedily admit cells with the highest utility-to-cost ratio until the budget $B$ is met.
-This policy can be deterministically computed by all agents, making it practically compatible with both centralized (V2I) and decentralized (V2V) operation (Sec. 5).
+This policy can be deterministically computed by all agents, making it practically compatible with both centralized (V2I) and decentralized (V2V) operation (Sec.
+5).
+We next show that this policy is optimal for the learned utility proxy–i.e.
+it maximizes $U$ among all feasible schedules, as formalized in Theorem 3.1 and 3.2.
 Figure 3: Visualization of meta utility map examples. The top-1 aggregated utility map eliminates redundancy by selecting the agent with the highest utility at each location.
-Train-Time Differentiable Scheduling. To address C2, we relax the discrete selection during training. Instead, we model the mask $m_{i}^{l}$ as the composition of two gates:
+Theorem 3.1(Singleton optimality)
+For each cell $l$, under the utility proxy in Eq. (4), any maximizer of $U^{l}$ selects at most one agent: $|\{i:m_{i}^{l}=1\}|\leq 1$.
+Proof
+Order agents by $u_{1}^{l}\geq u_{2}^{l}\geq\cdots$.
+Selecting agent $1$ yields $U^{l}=u_{1}^{l}$.
+Adding agent $2$ changes utility by $u_{2}^{l}-\min(u_{1}^{l},u_{2}^{l})=0$.
+Adding any $k\geq 3$ to a set $\mathcal{S}$ that already contains agent $1$ and $2$ yields marginal gain $u_{k}^{l}-\sum_{j\in\mathcal{S}}\min(u_{k}^{l},u_{j}^{l})\leq u_{k}^{l}-\min(u_{k}^{l},u_{2}^{l})=0.$
+Theorem 3.1 guarantees at most one agent per cell, inherently eliminating cross-agent redundancy (visualized in Fig.
+3).
+This reduces the joint agent-and-cell selection problem (C3) to selecting a subset of cells, each with a single pre-determined agent.
+Because every cell carries a $C$-channel feature vector plus an index, the per-cell cost is uniform for non-zero features [hu2022where2comm], and the resulting uniform-cost knapsack is optimally solved by selecting the cells with the largest utilities [kellerer2004multidimensional], as formalized in Theorem 3.2.
+Theorem 3.2(Greedy optimality under equal costs)
+Suppose each feature $(i,l)$ has the same nonzero cost $C_{i}^{l}\equiv c>0$.
+Let the per-frame budget be $B$ and define the integer capacity $K\triangleq\big\lfloor B/c\big\rfloor$.
+Then an optimal solution is obtained by selecting any $K$ features with the largest utilities $u_{i}^{l}$.
+We provide the proof to Theorem 3.2 in Appendix Sec.
+0.A.
+We illustrate the incorporation of the transmission policy into the end-to-end training pipeline below, which is essential in aligning the FUE-learned utility representation with the CP accuracy objective.
+Train-Time Differentiable Scheduling.
+Addressing the non-differentiable scheduling challenge (C2), we relax the discrete selection during training.
+Instead, we model the mask $m_{i}^{l}$ as the composition of two gates:
 (1) Importance gate (scalar, per agent–cell).
@@
 (2) Cross-agent top-1 (categorical, per cell).
-To model the “pick one agent per cell” constraint differentiably, we apply the Gumbel–Softmax reparameterization [jang2016categorical] on the vector of scores $\mathbf{u}^{l}=(u_{1}^{l},\dots,u_{N}^{l})$ with an independent Gumbel noise per agent: $g_{i}^{l}\sim\mathrm{Gumbel}(0,1)$ with $g_{i}^{l}=-\log(-\log\varepsilon_{i}^{l})$, $\varepsilon_{i}^{l}\sim\mathrm{Uniform}(0,1)$ , and a temperature $\gamma>0$, resulting in:
+To model the “pick one agent per cell” constraint differentiably, we apply the Gumbel–Softmax reparameterization [jang2016categorical] to the score vector $\mathbf{u}^{l}=(u_{1}^{l},\dots,u_{N}^{l})$ with independent Gumbel noise $g_{i}^{l}\sim\mathrm{Gumbel}(0,1)$ and temperature $\gamma>0$:
 $\boldsymbol{\beta}_{i}^{l}(\gamma)=\mathrm{softmax}_{i}\left(\frac{u_{i}^{l}+g_{i}^{l}}{\gamma}\right)=\frac{\exp\left((u_{i}^{l}+g_{i}^{l})/\gamma\right)}{\sum_{k=1}^{N}\exp\left((u_{k}^{l}+g_{k}^{l})/\gamma\right)}.$ (5)
@@
 In the backward pass, we compute gradients as if the mask was the soft, differentiable approximation $\tilde{m}_{i}^{l}=\alpha_{i}^{l}\boldsymbol{\beta}_{i}^{l}(\gamma)$.
-$\gamma$ follows the same annealing schedule as $\eta$ during training.
+$\gamma$ follows the same annealing schedule as $\eta$.
 Inference-time Deterministic Policy.
@@
 We sort all $(i,l)$ with $m_{i}^{l}=1$ by $\rho_{i}^{l}$ and admit the longest prefix whose cumulative cost $\leq B$.
-Proposition 1 (Consistency of differentiable and deterministic schedulers).
+Proposition 1(Consistency of differentiable and deterministic schedulers)
 For any $(i,l)$ with $u_{i}^{l}\neq\tau$ and with no injected noise at inference, we have
@@
 $\lim_{\eta\to 0^{+}}\tilde{m}_{i}^{l}=\mathbb{I}[u_{i}^{l}=\max_{k}u_{k}^{l}]\cdot\mathbb{I}[u_{i}^{l}>\tau]=m_{i}^{l}.$ (10)
-Prop.
-(1) indicates that the train-time relaxed mask converges pointwise to the deterministic policy.
-Also, instead of a heuristic design, we show that this policy is provably optimal for the utility proxy in Eq.
-(4) and the knapsack problem under uniform costs.
-Theorem 1 (Singleton optimality).
-For each cell $l$, under the utility proxy in Eq. (4), any maximizer of $U^{l}$ selects at most one agent: $|\{i:m_{i}^{l}=1\}|\leq 1$.
-Proof.
-Order agents by $u_{1}^{l}\geq u_{2}^{l}\geq\cdots$.
-Selecting agent $1$ yields $U^{l}=u_{1}^{l}$.
-Adding agent $2$ changes utility by $u_{2}^{l}-\min(u_{1}^{l},u_{2}^{l})=0$.
-Adding any $k\geq 3$ to a set $\mathcal{S}$ that already contains agent $1$ and $2$ yields marginal gain $u_{k}^{l}-\sum_{j\in\mathcal{S}}\min(u_{k}^{l},u_{j}^{l})\leq u_{k}^{l}-\min(u_{k}^{l},u_{2}^{l})=0.$
-Figure 4: Visualization of feature map examples.
-Theorem (1) ensures that there is at most one agent with the highest utility to transmit per location cell, inherently eliminating cross-agent redundancy, as visualized in Fig.
-3.
-This reduces the knapsack problem (C3) of selecting features across both agent and location dimensions to selecting the best set of location cells to transmit.
-Although the problem remains a knapsack problem, it becomes a uniform-cost knapsack, since the number of channels is fixed per feature, resulting in the same transmission cost.
-It is known that the uniform-cost knapsack reduces to a cardinality-constrained selection, which can be optimally solved by taking the elements with the largest utilities [kellerer2004multidimensional], as shown in Theorem (2).
-Theorem 2 (Greedy optimality under equal costs).
-Suppose each feature $(i,l)$ has the same nonzero cost $C_{i}^{l}\equiv c>0$.
-Let the per-frame budget be $B$ and define the integer capacity $K\triangleq\big\lfloor B/c\big\rfloor$.
-Then an optimal solution is obtained by selecting any $K$ features with the largest utilities $u_{i}^{l}$.
-Proof.
-By Theorem (1), each cell yields at most one admissible candidate; thus the global problem is
-$\max_{\mathcal{S}\subseteq\mathcal{C}}\ \sum_{(i,l)\in\mathcal{S}}u_{i}^{l}\quad\text{s.t. }\ |\mathcal{S}|\leq K,$ (11)
-where $\mathcal{C}$ is the set of at-most-one-per-cell candidates and all costs equal $c$.
+Prop. 1 indicates that the train-time policy converges pointwise to the inference policy.
+Figure 4: Feature map visualization. The semantic loss and learnable threshold $\kappa$ produce sparse features that concentrate activations on semantically relevant regions.
+3.5 End-to-End Communication-Aware Training
+We unify these components into a single end-to-end trainable framework. The final loss combines the downstream task loss (e.g. focal loss) with the semantic loss:
+$\mathcal{L}_{\text{JigsawComm}}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{semantic}},$ (11)
+where $\lambda$ balances feature sparsity and task accuracy.
+The $\mathcal{L}_{\text{task}}$ gradient flows through the differentiable scheduler (Sec.
+3.4) to train the FUE head (Sec.
+3.3), while $\mathcal{L}_{\text{semantic}}$ shapes the encoder (Sec.
+3.2).
+By jointly learning the sparsity threshold $\kappa$ and utility threshold $\tau$, JigsawComm activates and transmits features only when they contribute to CP accuracy, as illustrated in Fig.
+4.
+We note that a naive alternative—penalizing the transmission masks $\{\mathbf{M}_{i}\}$ directly—performs worse in practice because it penalizes all communication indiscriminately, disregarding semantic importance and redundancy.
+3.6 Robustness Enhancement
+Robustness to Misalignment.
+The top-1-per-cell policy in Eq.
+(7) compares utilities $u_{i}^{l}$ across agents at each spatial cell $l$.
+In practice, localization noise and communication delays introduce translational offsets between agents’ coordinate frames [zhang2023robust], causing the scheduler to compare misaligned cells.
+To address this, we introduce a lightweight alignment step that estimates and corrects offsets using only the already-exchanged meta utility maps $\{\mathbf{U}_{i}\}$, requiring no additional communication.
+Inspired by the feature matching and warping in [Zhang_2024_CVPR], for each non-ego agent $j$, we extract the top-$K$ cells by utility value from both $\mathbf{U}_{\text{ego}}$ and $\mathbf{U}_{j}$ as sparse keypoint sets and estimate a translational offset $\boldsymbol{\delta}_{j}$ via RANSAC [derpanis2010ransac]: in each iteration, we sample one point from each set, compute the candidate offset, and count inliers within $\varepsilon$ cells.
+The best hypothesis is refined by averaging inlier correspondences.
+Offsets with low confidence are discarded to avoid interpolation artifacts.
+Given the estimated offsets, each non-ego utility map $\mathbf{U}_{j}$ and features $\mathbf{F}_{j}$ are warped by $\boldsymbol{\delta}_{j}$ via bilinear grid sampling before scheduling and fusion.
+Robustness to Packet Loss.
+The top-1-per-cell policy also provides a natural fallback under packet loss: if a non-ego agent’s scheduled feature is dropped during transmission, the ego vehicle uses its own local feature at the affected cell, requiring no additional communication.
+Because the ego’s feature is always available and semantically meaningful in visible areas, this fallback degrades gracefully (Appendix 0.E) and guarantees the performance lower-bounded by local detection.
+Alternatively, a Top-$K$ policy could pre-select backup agents when more bandwidth is available, where the scheduler pre-selects $K$ candidate agents per cell and falls back to the next-best candidate upon packet loss.
+4 Evaluation
+Table 1: Main results on OPV2V and DAIR-V2X.
+All methods transmit their full output (unlimited budget $B$).
+Rel.
+Eff.
+is the mAP@0.5-to-total-size ratio, normalized to JigsawComm ($100\%$).
+JigsawComm achieves SOTA accuracy while reducing data volume by over $500{\times}$.
+OPV2V DAIR-V2X
+Method mAP@.5 mAP@.7 Per-Ag. Total Rel. Eff. mAP@.5 mAP@.7 Per-Ag. Total Rel. Eff.
+(KB) (KB) (%) (KB) (KB) (%)
+No Fusion 0.68 0.60 0.50 0.44
+F-Cooper [f_cooper] 0.89 0.79 8,800 22,792 0.19 0.56 0.39 16,384 32,768 0.70
+V2X-ViT [xu2022v2x] 0.87 0.67 8,800 22,792 0.18 0.68 0.50 16,384 32,768 0.84
+[…]
+$\ast$ + Enc-Dec 0.89 0.81 4,400 11,396 0.38 0.65 0.51 8,192 16,384 1.61
+Where2comm [hu2022where2comm] 0.85 0.60 1,152 2,795 1.47 0.59 0.46 6,295 12,590 1.91
+ERMVP [Zhang_2024_CVPR] 0.89 0.79 845 2,189 1.97 0.64 0.51 2,914 5,827 4.47
+CoST [tang2025cost] 0.87 0.76 4,194 12,163 0.35 0.62 0.47 4,194 8,389 3.01
+[…]
+4.1 Experimental Setup
+Datasets.
+We evaluate JigsawComm on two CP benchmarks: (i) OPV2V [xu2022opv2v], a large-scale dataset with 11,464 LiDAR frames from 2–5 agents across 70 scenes in 8 CARLA [Dosovitskiy17Carla] towns; and (ii) DAIR-V2X [yu2022dair], the first large-scale real-world CP dataset for vehicle-infrastructure cooperation, containing 71,254 frames.
+Implementation Details.
+We adopt PointPillars [lang2019pointpillars] as the LiDAR BEV backbone and a max-out module as the fusion backbone [f_cooper].
+We train JigsawComm for up to 100 epochs using Adam ($\beta_{1}\!=\!0.9,\beta_{2}\!=\!0.999$), with a base learning rate of 0.002.
+The Gumbel-Softmax temperatures $\eta,\gamma$ are annealed from $0.9\to 0.1$.
+Evaluation Metrics.
+We evaluate 3D object detection performance using the mean Average Precision (mAP) at IoU thresholds of 0.5 and 0.7.
+Communication cost is measured by the Average Total Data Size (KB) transmitted per frame by all agents, under FP8.
+We also report Relative Communication Efficiency, defined as $\text{mAP}@0.5/\text{Total Size}$, normalized to our method (100%) for comparison.
+The required bandwidth is calculated as the total transmitted feature size multiplied by the frame rate ($r=10$ FPS).
+Baselines.
+We follow the standard evaluation protocol in communication-efficient CP [xu2025cosdh, hu2024communication, mao2025diffcp, zhao2025quantv2x] and evaluate: (i) a no-fusion single-vehicle detector; (ii) full-transmission intermediate-fusion (F-Cooper [f_cooper], V2X-ViT [xu2022v2x], AttFusion [xu2022opv2v]); and (iii) communication-efficient methods, including the canonical baseline Where2comm [hu2022where2comm] and prior art ERMVP (CVPR’24 [Zhang_2024_CVPR]) and CoST (ICCV’25 [tang2025cost]).
+All baselines are reproduced using their official codebases with the PointPillars backbone [lang2019pointpillars] for a fair comparison.
+4.2 Quantitative Results
+Main Results.
+Table 1 presents our main comparison.
+On OPV2V, JigsawComm matches or exceeds all baselines in accuracy, including the best full-transmission method AttFusion, while using only 44 KB per frame—a $520{\times}$ reduction from AttFusion’s 22,792 KB.
+On the real-world DAIR-V2X dataset, JigsawComm achieves the highest mAP@0.5 (0.69), surpassing V2X-ViT (0.68) while requiring only 282 KB ($116{\times}$ less data).
+Among communication-efficient baselines, JigsawComm outperforms the strongest competitor ERMVP by 3 mAP points on OPV2V and 5 on DAIR-V2X, while using $50{\times}$ and $21{\times}$ less bandwidth, respectively, thanks to the joint encoder–scheduler design eliminating both intra- and inter-agent redundancy simultaneously.
+(a) OPV2V
+(b) DAIR-V2X
+Figure 5: Accuracy vs. required bandwidth. JigsawComm (curve) is flexible via $B$; baselines (points) are fixed.
+(a) Accuracy
+(b) Feature size
+Figure 6: Scalability on OPV2V by #agents. JigsawComm bounds the payload to $\mathcal{O}(1)$.
+Accuracy–Bandwidth Tradeoff.
+Fig.
+5 plots JigsawComm’s accuracy as a function of the available bandwidth.
+On OPV2V, JigsawComm reaches 0.90 mAP@0.5 at only $\sim$}200 Kbps and saturates beyond $\sim$}1,000 Kbps, well within the V2X capacity ($\leq$}20 Mbps)[FCC_DSRC, CV2X_ETSI_EN_303_613].
+On DAIR-V2X, the curve saturates at a similarly low bandwidth, confirming that the learned utility proxy successfully identifies the small fraction of features that drive perception accuracy.
+This property is especially valuable for congested V2X channels, where the scheduler can gracefully degrade by transmitting only the highest-utility cells first.
+Note that a single trained model serves the entire curve; the budget $B$ acts as a simple “knob” at inference.
+Scalability with Number of Agents.
+Fig.
+6 validates $\mathcal{O}(1)$ data-channel scaling on OPV2V.
+The total feature size grows only slightly from $\approx$}40 KB (N=2N{=}2) to $\approx$}58 KB (N=4N{=}4) due to increased coverage, while accuracy remains high.
+This sub-linear growth is a direct consequence of the top-1-per-cell policy (Theorem 3.1): as more agents observe the same region, the scheduler selects only the single best contributor, converting redundancy into bandwidth savings.
+Note that the OPV2V benchmark groups scenes by the number of participating agents rather than incrementally adding agents to a fixed scene.
+Consequently, differences in scene composition (object density, spatial layout) across groups can cause slight non-monotonicity in the per-group averages (e.g.
+a small decrease from N=4N{=}4 to N=5N{=}5).
+The overall sub-linear trend nevertheless holds clearly across the range.
+We further test on a custom dataset with 10–20 CAVs (Appendix 0.D): the payload saturates beyond $\sim$}15 agents, confirming $\mathcal{O}(1)$ scaling as inter-agent overlap dominates.
+End-to-End Efficiency.
+The average end-to-end latency is $45.82{\pm}26.96$ ms, composed of $28.27{\pm}23.88$ ms computation on a single RTX 4090 and $17.55{\pm}7.14$ ms communication at 20 Mbps, which is well within the 100 ms real-time budget for autonomous driving [ADConstraints].
+The meta utility map overhead from all agents is only $0.08{\pm}0.04$ KB per frame, an average of less than $0.03$ KB/agent under FP4.
+4.3 Robustness Analysis
+Real-world V2X deployments are subject to imperfect localization and communication delays.
+We evaluate JigsawComm’s robustness to both factors on the OPV2V dataset and compare against ERMVP [Zhang_2024_CVPR] and CoST [tang2025cost].
+We further show JigsawComm’s robustness against packet drops in Appendix 0.E.
+(a) Localization error
+(b) Delay
+Figure 7: Robustness on OPV2V. JigsawComm degrades gracefully under localization noise and communication delay.
+(a) OPV2V
+(b) DAIR-V2X
+Figure 8: Top-1 vs. Top-2 agents per cell. Redundant features degrade accuracy, corroborating Theorem 3.1.
+Localization Error.
+We inject Gaussian pose noise with $\sigma\in\{0,0.2,\ldots,1.0\}$ m into each agent’s reported position.
+As shown in Fig.
+7(a), JigsawComm degrades gracefully from 0.90 ($\sigma{=}0$) to 0.74 ($\sigma{=}1.0$ m), a relative drop of 18%, thanks to the RANSAC-based utility map alignment (Sec.
+3.6).
+In contrast, CoST drops from 0.87 to 0.39 (55%).
+The resilience of JigsawComm stems from the fact that the sparse, high-utility cells used for RANSAC matching are predominantly located on or near objects, providing robust correspondences even under noise.
+At the practically relevant range of $\sigma{\leq}0.4$ m, JigsawComm’s accuracy remains above 0.85.
+Communication Delay.
+We simulate delays of $\{0,100,\ldots,500\}$ ms by substituting each agent’s current features with those from an earlier timestamp.
+Fig.
+7(b) shows JigsawComm degrades more gracefully: at 200 ms it retains AP@0.5 of 0.72 (vs.
+0.59 ERMVP, 0.41 CoST); at 500 ms—equivalent to 5 missed frames at 10 FPS—JigsawComm still achieves 0.63, outperforming both baselines by a large margin.
+We attribute this resilience to two complementary effects: (i) the scheduler selects only the most informative features, so the fused representation is less sensitive to the staleness of any individual contribution; and (ii) the meta utility map and feature warping via RANSAC compensate for spatial misalignment induced by asynchronous transmission [zhang2023robust].
+4.4 Ablation Study
+Table 2: Ablation study of our core components.
+OPV2V DAIR-V2X
+$\mathcal{L}_{\text{sem}}$ FUE mAP@0.5 Size (KB) Rel. Eff. (%) mAP@0.5 Size (KB) Rel. Eff. (%)
+0.93 26,292.4 0.17 0.55 22,670.6 0.99
+0.94 13,516.8 0.33 0.61 7,375.0 3.38
+0.92 43.8 100 0.69 282.0 100
+Impact of Core Components.
+Table 2 ablates $\mathcal{L}_{\text{semantic}}$ and the FUE & Scheduler.
+Using $\mathcal{L}_{\text{semantic}}$ alone produces sparse features but transmits them in full, yielding 26,292 KB on OPV2V because inter-agent redundancy is unaddressed.
+Using the FUE alone selects features intelligently but operates on dense representations (13,517 KB).
+Only when both are combined does transmission drop to 44 KB while maintaining high accuracy: $\mathcal{L}_{\text{semantic}}$ ensures each cell carries only task-relevant content (intra-agent), while the scheduler selects only the best contributor per cell (inter-agent).
+On DAIR-V2X, the synergy is even more pronounced: the full system achieves 0.69 mAP@0.5—substantially higher than either component alone (0.55 or 0.61)—because the real-world dataset’s viewpoint diversity makes both forms of redundancy elimination jointly beneficial.
+Effect of Redundancy.
+Theorem 3.1 predicts that selecting more than the Top-1 agent per cell is suboptimal under our utility proxy.
+Fig.
+8 validates this empirically: a Top-2 policy that transmits features from the two highest-utility agents per cell is worse than Top-1 at nearly every bandwidth point on both datasets.
+This “more is worse” effect stems from a limitation of current fusion algorithms dominantly adopted by SOTA CP algorithms (e.g.
+max-pooling, attention [xu2022opv2v, hu2022where2comm, f_cooper]).
+The fused feature can be expressed as a weighted sum of the features from different agents.
+If the second-best agent’s feature is noisier or semantically inconsistent, it dilutes the clean, high-utility representation rather than reinforcing it.
+This corroborates JigsawComm’s design of selecting only the highest-quality, non-redundant features.
+5 Discussion
+Practical V2X Deployment.
+JigsawComm maps naturally onto the established dual-channel V2X architecture [FCC_DSRC, CV2X_ETSI_EN_303_613]: meta utility maps on the control channel, sparse BEV features on the data channel via the Collective Perception Message [cpm].
+The scheduler’s utility-ranked priorities integrate directly into both scheduling-based (e.g.
+C-V2X [CV2X_ETSI_EN_303_613]) and contention-based protocols (e.g.
+DSRC [FCC_DSRC]), supporting centralized (V2I) and decentralized (V2V) modes for participants in the same communication range.
+Limitations.
+The optimality transmission guarantees (Theorems 3.13.2) hold under the learned utility proxy as a data-driven intermediate representation, not the true task-level objective (which is unavailable during inference for optimizing the transmission decision); their practical value thus depends on how well the proxy tracks perception accuracy (validated empirically in Sec.
+4).
+The current design assumes a shared BEV backbone across agents; heterogeneous architectures would require feature alignment before utility comparison [lu2024heal].
+Finally, the “more is worse” finding (Sec.
+4.4) highlights a limitation of existing fusion modules [f_cooper, xu2022opv2v, xu2022v2x, hu2022where2comm]; developing fusion architectures that can leverage multi-agent redundancy synergistically remains an open problem.
+6 Conclusion and Future Work
+[…]
+Future Work.
+Several promising directions remain.
+First, the sparse features produced by JigsawComm are amenable to entropy coding, which could further reduce the transmitted payload.
+Second, extending the framework to joint source-channel coding would enable end-to-end optimization over unreliable wireless channels rather than treating compression and transmission separately.
+Third, adapting the framework to camera-based BEV features and multi-modal (LiDAR + camera) fusion would broaden its applicability to the heterogeneous sensor suites.
+Supplementary Material
+Appendix 0.A Proof to Theorem 3.2
+Proof
+By Theorem 3.1, each cell yields at most one admissible candidate; thus the global problem is $\max_{\mathcal{S}\subseteq\mathcal{C}}\ \sum_{(i,l)\in\mathcal{S}}u_{i}^{l}\;\text{s.t.
+}\ |\mathcal{S}|\leq K,$ where $\mathcal{C}$ is the set of at-most-one-per-cell candidates and all costs equal $c$.
 This is a cardinality-constrained selection problem.
@@
 take $|\mathcal{O}|=K$.
-Order items in $\mathcal{G}$ and $\mathcal{O}$ so that $u(g_{1})\geq\cdots\geq u(g_{K})$ and $u(o_{1})\geq\cdots\geq u(o_{K})$.
-Since $\mathcal{G}$ contains the top-$K$ utilities overall, we must have $u(g_{j})\geq u(o_{j})$ for every $j=1,\dots,K$.
-Summing over $j$ yields $\sum_{j=1}^{K}u(g_{j})\geq\sum_{j=1}^{K}u(o_{j})$, contradicting $\sum_{o\in\mathcal{O}}u(o)>\sum_{g\in\mathcal{G}}u(g)$.
+Order items so that $u(g_{1})\geq\cdots\geq u(g_{K})$ and $u(o_{1})\geq\cdots\geq u(o_{K})$.
+Since $\mathcal{G}$ contains the top-$K$ utilities overall, $u(g_{j})\geq u(o_{j})$ for every $j=1,\dots,K$.
+Summing yields $\sum_{j}u(g_{j})\geq\sum_{j}u(o_{j})$, a contradiction.
 Hence $\mathcal{G}$ is optimal.
-∎
-3.5 End-to-End Communication-Aware Training
-We unify these components into an end-to-end trainable framework. The final loss combines the downstream task loss (e.g. focal, cross-entropy, etc.) with the semantic loss:
-$\mathcal{L}_{\text{JigsawComm}}=\mathcal{L}_{\text{task}}+\lambda\mathcal{L}_{\text{semantic}},$ (12)
-where $\lambda$ is a hyperparameter balancing the sparsity of the encoded feature and task accuracy.
-The $\mathcal{L}_{\text{task}}$ gradient flows through the differentiable scheduler (Sec.
-3.4) to train the FUE head (Sec.
-3.3), while $\mathcal{L}_{\text{semantic}}$ shapes the encoder (Sec.
-3.2).
-By learning the sparsity threshold $\kappa$ and the utility threshold $\tau$ jointly, JigsawComm learns to activate and transmit features only when they contribute to CP accuracy, maximizing the utility of every transmitted bit, as shown in Fig.
-4.
-4 Evaluation
+Table 3: Detailed architectural specifications for JigsawComm.
+NN is the number of cooperative agents in a scene, $P$ the number of non-empty pillars, and $A$ the number of anchors per cell.
+With voxel resolution $0.4$ m and the default LiDAR range, the initial BEV grid is $H_{0}{\times}W_{0}{=}200{\times}704$; after the multi-scale backbone the feature map is $H{\times}W{=}100{\times}352$.
+Output size JigsawComm framework
 […]
-Dataset OPV2V DAIR-V2X
-Model mAP @0.5 mAP @0.7 Per-Agent Size (KB) Total Data Size (KB) Rel. Eff. (%) mAP @0.5 mAP @0.7 Per-Agent Size (KB) Total Data Size (KB) Rel. Eff. (%)
-No Fusion 0.68 0.60 - - - 0.50 0.44 - - -
-F-Cooper 0.89 0.79 8,800.00 22,792.00 0.19 0.56 0.39 16,384.00 32,768.00 0.70
-V2X-ViT 0.87 0.67 8,800.00 22,792.00 0.18 0.68 0.50 16,384.00 32,768.00 0.84
-AttFusion()(*) 0.91 0.82 8,800.00 22,792.00 0.19 0.65 0.51 16,384.00 32,768.00 0.81
-()(*) + Enc-Dec 0.89 0.81 4,400.00 11396.00 0.38 0.65 0.51 8,192.00 16,384.00 1.61
-Where2comm 0.85 0.60 1,152.47 2,794.72 1.47 0.46 0.23 14,692.88 29,385.76 0.64
-JigsawComm 0.92 0.82 17.46 43.84 100.00 0.69 0.52 148.96 281.97 100.00
-4.1 Experimental Setup
-Dataset.
-We evaluate JigsawComm on two CP datasets: (i) OPV2V [xu2022opv2v], which is a large-scale benchmark collected across 70 diverse scenes from 8 towns in the digital twin simulator CARLA [Dosovitskiy17Carla].
-It includes 11,464 frames of LiDAR point cloud data from 2-5 agents and 232,913 annotated 3D bounding boxes.
-(ii) DAIR-V2X [yu2022dair], which is the first large-scale real-world CP dataset for vehicle-infrastructure cooperation, consisted of 71,254 frames.
-Implementation Details.
-We adopt PointPillars [lang2019pointpillars] as the LiDAR BEV backbone and a max-out module as the fusion backbone [f_cooper].
-We train JigsawComm for up to 100 epochs using Adam ($\beta_{1}\!=\!0.9,\beta_{2}\!=\!0.999$), with a base learning rate of 0.002.
-The Gumbel-Softmax temperatures $\eta,\gamma$ are annealed from $0.9\to 0.1$.
+Pillar Scatter $N\times 64\times H_{0}\times W_{0}$ Scatter pillar features to BEV grid
+[…]
+[…]
+[…]
+BEV Backbone (decoder) $N\times 384\times H\times W$
+[…]
+[…]
+[…]
+Concat3, 384
+Sparse Thresholding $N\times 384\times H\times W$ $\mathbf{F}_{i}\leftarrow\mathbf{F}_{i}\odot\mathbb{I}[\mathbf{F}_{i}>\kappa]$ ,   $\kappa$ learnable
+[…]
+Transmission Scheduler $N\times 1\times H\times W$
+Train:  $\tilde{m}_{i}^{l}=\sigma\!\big(\tfrac{u_{i}^{l}-\tau}{\eta}\big)\cdot\boldsymbol{\beta}_{i}^{l}(\gamma)$(STE)
+Infer:  $m_{i}^{l}=\mathbb{I}[i{=}\arg\max_{k}u_{k}^{l}]\cdot\mathbb{I}[u_{i}^{l}{\geq}\tau]$
+Budget:  greedy top-$K$ by $u_{i}^{l}$
+Fusion $1\times 384\times H\times W$ Element-wise max over $N$ agents
+Detection Head $H\times W\times A$
+[…]
+[…]
+Appendix 0.B Architectural Configurations
+Given the definitions in Sec. 3, the entire JigsawComm pipeline can be concisely formulated as follows:
+$\mathbf{z}_{i}$ $=\mathrm{PillarVFE}(X_{i}),$ $\mathbf{z}_{i}$ $\in\mathbb{R}^{P_{i}\times 64},$ $\text{for agent }i$ (12)
+$\mathbf{F}_{i}$ $=\Phi_{\text{enc}}\!\big(\mathrm{Scatter}(\mathbf{z}_{i})\big),$ $\mathbf{F}_{i}$ $\in\mathbb{R}^{C\times H\times W},$ (13)
+$\mathbf{F}_{i}$ $\leftarrow\mathbf{F}_{i}\odot\mathbb{I}[\mathbf{F}_{i}>\kappa],$ (14)
+$u_{i}^{l}$ $=\mathrm{ReLU}\!\big(\mathbf{w}^{\top}\mathbf{f}_{i}^{l}+b\big),$ $\mathbf{U}_{i}$ $\in\mathbb{R}^{1\times H\times W},$ (15)
+$m_{i}^{l}$ $=\mathrm{Sched}\!\big(\{u_{j}^{l}\}_{j=1}^{N},\,\tau,\,B\big),$ $\mathbf{M}_{i}$ $\in\{0,1\}^{H\times W},$ (16)
+$\hat{Y}$ $=\Phi_{\text{dec}}\!\Big(\Phi_{\text{fuse}}\!\big(\{\mathbf{M}_{i}\odot\mathbf{F}_{i}\}_{i=1}^{N}\big)\Big),$ (17)
+where $X_{i}$ denotes the raw LiDAR point cloud captured by agent $a_{i}$.
+These points are voxelized and fed into the Pillar VFE [lang2019pointpillars], yielding $P_{i}$ pillar features $\mathbf{z}_{i}$ of dimension $64$.
+The pillars are scattered onto a 2D BEV grid of size $H_{0}\times W_{0}$ and processed by the multi-scale BEV backbone encoder $\Phi_{\text{enc}}$, which consists of three downsampling blocks with $3$, 55, and $8$ convolutional layers at $64$, $128$, and $256$ channels respectively (each with stride-2 entry, BatchNorm, and ReLU), followed by three transposed-convolution upsampling blocks whose outputs are concatenated to produce C=384C{=}384-channel features $\mathbf{F}_{i}$ at half the input spatial resolution ($H{=}H_{0}/2$, $W{=}W_{0}/2$).
+The sparse thresholding (Eq.
+14) zeros out activations below the learnable threshold $\kappa$, producing compact features with exact structural sparsity.
+The FUE head (Eq.
+15) is a single $1{\times}1$ convolution that maps each $C$-dimensional feature cell to a scalar utility $u_{i}^{l}$, forming the meta utility map $\mathbf{U}_{i}$.
+After exchanging the sparse utility maps among agents, the transmission scheduler (Eq.
+16) deterministically computes the binary mask $\mathbf{M}_{i}$ via the per-cell top-1 selection, threshold $\tau$, and budget constraint $B$ (Sec.
+3.4).
+The masked features $\mathbf{M}_{i}\odot\mathbf{F}_{i}$ are fused via element-wise max pooling across agents, and the fused representation is passed through two $1{\times}1$ convolution heads for classification ($A$ anchors) and bounding-box regression ($7A$ outputs).
+The detailed layer-by-layer specifications are given in Table 3.
+Notably, JigsawComm introduces two lightweight additions beyond the standard PointPillar backbone [lang2019pointpillars]: (i) the FUE head (C+1=385C{+}1=385 parameters) and (ii) two learnable scalar thresholds $\kappa$ and $\tau$.
+The transmission scheduler is purely algorithmic with no learnable parameters.
+The additional parameter overhead is therefore negligible ($0.01\%$ of total model parameters).
+Appendix 0.C Training Procedure
+Optimization.
+We use Adam [kingma2014adam] with initial learning rate $2{\times}10^{-3}$, weight decay $10^{-4}$, and $\epsilon{=}10^{-10}$.
+The learning rate is decayed by a factor of $0.1$ at epochs $15$ and $30$ using a multi-step schedule.
+All models are trained for $50$ epochs with batch size $4$ on a single NVIDIA RTX 4090 GPU.
+We save checkpoints every $2$ epochs and select the model with the lowest validation loss.
 We select $\lambda=1000$ for OPV2V and $\lambda=100$ for DAIR-V2X, through standard cross validation.
-Evaluation Metrics.
-We evaluate 3D object detection performance using the mean Average Precision (mAP) at IoU thresholds of 0.5 and 0.7.
-Communication cost is measured by the Average Total Data Size (KB) transmitted per frame by all agents, under FP8.
-We also report Relative Communication Efficiency, defined as $\text{mAP}@0.5/\text{Total Size}$, normalized to our method (100%) for comparison.
-The required bandwidth is calculated as the total transmitted feature size multiplied by the frame rate ($r=10$ FPS).
-Baselines.
-We follow the evaluation protocol that has become standard in communication-efficient CP [xu2025cosdh, hu2024communication, mao2025diffcp, zhao2025quantv2x]: (i) a no-fusion single-vehicle detector, (ii) strong full-transmission intermediate-fusion methods that transmit dense BEV features, including F-Cooper [f_cooper], V2X-ViT [xu2022v2x], and AttFusion [xu2022opv2v], and (iii) communication-efficient CP via Where2comm [hu2022where2comm], which is the de facto canonical communication-efficient baseline adopted by recent works [xu2025cosdh, hu2024communication, mao2025diffcp, zhao2025quantv2x].
-Although some recent approaches [hu2024communication, mao2025diffcp, zhao2025quantv2x] achieve comparable bandwidth requirements, they suffer from practical limitations in generalizability or computation latency, as discussed in Sec.
-2.2.
-4.2 Quantitative Results
-[…]
-(a) OPV2V
-(b) DAIR-V2X
-Figure 5: Tradeoff between accuracy and required bandwidth.
-Accuracy-Bandwidth Tradeoff.
-Figure 5 visualizes JigsawComm’s performance versus the required communication bandwidth ($BW$).
-The baselines are fixed points, representing inflexible, inefficient strategies.
-JigsawComm’s curve is generated by varying the communication budget $B$.
-The accuracy of JigsawComm saturates quickly after $BW\geq$1,000 Kbps, making it a flexible, generalizable, and pragmatic solution for deployment under the limited bandwidth offered by V2X (e.g, $\leq$20 Mbps) [FCC_DSRC, CV2X_ETSI_EN_303_613].
-Figure 6: Accuracy, feature size, and required bandwidth grouped by agent counts on OPV2V.
-Instead of a linear increase in total feature size with more agents, JigsawComm induces an asymptotic $\mathcal{O}(1)$ communication cost while achieving high accuracy.
-Scalability with Number of Agents.
-JigsawComm achieves an asymptotic $\mathcal{O}(1)$ communication cost, where the total bandwidth is bounded by the scene’s information content in a single sparse feature, not the number of participating agents $N$.
-We validate this claim empirically in Fig.
-6 using the OPV2V dataset, which includes 2 to 5 agents.
-JigsawComm’s accuracy remains high as the number of agents varies.
-The total feature size grows slightly from $\approx$40 KB at $N=2$ to $\approx$58 KB at $N=4$, which is attributed to the fact that more agents increases collective scene coverage.
-Our FUE correctly identifies these non-redundant cells as having high utility and adds them to the ‘jigsaw puzzle’.
+Temperature Annealing.
+The Gumbel–Softmax temperature $\gamma$ and importance-gate temperature $\eta$ share the same annealing schedule: both are initialized to $0.9$ and multiplied by a decay factor of $0.9$ each epoch, yielding $\eta_{t}=0.9^{t+1}$ at epoch $t$.
+The temperature is clamped at a minimum of $0.01$ to maintain numerical stability.
+As $\eta\to 0$, the differentiable masks converge to the deterministic inference-time policy (Prop.
+1).
+Data Augmentation.
+During training, we apply three standard augmentations to the global point cloud: (i) random flip along the $x$-axis, (ii) random rotation uniformly sampled from $\pi/4,\pi/4]$, and (iii) random scaling with a factor uniformly sampled from $[0.95,1.05]$.
+Bandwidth Control.
+During training, no explicit bandwidth constraint is imposed ($B{=}\infty$); the semantic loss $\mathcal{L}_{\text{semantic}}$ and the learnable threshold $\tau$ naturally drive the model toward compact transmission.
+At inference, the budget $B$ can be set to any desired level: the scheduler simply admits cells by descending utility until the budget is met (Sec.
+3.4).
+This allows a single trained model to operate at any point on the bandwidth–accuracy trade-off curve without retraining.
 Figure 9: JigsawComm’s total feature size and required bandwidth under 10 FPS on our custom dataset.
-Since OPV2V is limited to 5 agents, we conduct further experiments on a custom-simulated dataset to analyze the asymptotic trend.
-We utilized the CARLA digital twin simulator [Dosovitskiy17Carla] in conjunction with the OpenCDA framework [opencda] for coordinated multi-agent simulation and data recording.
+$p_{\text{drop}}$ AP@.3 AP@.5 AP@.7
+0.00 0.933 0.925 0.826
+0.05 0.930 0.922 0.822
+0.10 0.926 0.919 0.817
+0.15 0.923 0.915 0.813
+0.20 0.919 0.911 0.808
+Table 4: Robustness to packet loss on OPV2V. $p_{\text{drop}}$: per-agent drop probability; ego fallback at affected cells.
+Appendix 0.D Scaling with More Agents
+To further test JigsawComm’s ability to scale beyond the 2-5 agents available from the OPV2V and DAIR-V2X dataset Sec.
+4.1, we conducted further experiments on a custom collected dataset.
+We utilized the CARLA digital driving simulator [Dosovitskiy17Carla] in conjunction with the OpenCDA framework [opencda] for coordinated multi-agent simulation and data recording.
+The data was collected following the same format and conventions as the public OPV2V benchmark dataset, including sensor configurations, coordinate systems, and CAV parameters, to ensure compatibility with existing CP models and evaluation pipelines.
 We set up the scene within CARLA’s ‘Town10HD’ map with 10–20 CAVs following realistic traffic rules, at a large multi-lane intersection.
-The data was collected following the same format and conventions as the public OPV2V benchmark dataset, including sensor configurations, coordinate systems, and CAV parameters, to ensure compatibility with existing CP models and evaluation pipelines.
 Fig.
 9 presents the result of JigsawComm on the custom dataset.
-We can observe that JigsawComm’s total communication cost does not double when scaling from 10 to 20 agents.
-It saturates and remains relatively flat when the number of agents is greater than 15, empirically verifying the asymptotic $\mathcal{O}(1)$ communication cost claim.
+We can observe that JigsawComm’s total data-channel payload does not double when scaling from 10 to 20 agents.
+It saturates and remains relatively flat when the number of agents is larger than 15, confirming the asymptotic $\mathcal{O}(1)$ data-channel scaling.
 As more agents join, the collective coverage saturates, the overlap becomes dominant, and the marginal contribution of each new agent becomes smaller.
-End-to-End Efficiency.
-Beyond theoretical bandwidth savings, we validate JigsawComm’s end-to-end efficiency.
-The average end-to-end latency is $45.82\pm 26.96$ ms, which is well within the stringent 100 ms real-time budget for autonomous driving perception [ADConstraints].
-This total latency consists of $28.27\pm 23.88$ ms for computation on an RTX 4090 GPU and $17.55\pm 7.14$ ms for communication (assuming a 20 Mbps bandwidth constraint), including the total feature payload from all agents and the meta utility maps with a small total overhead of $0.08\pm 0.04$KB.
-4.3 Ablation Study
-Table 3: Ablation study of our core components.
-Components OPV2V DAIR-V2X
-Semantic
-Loss
-FUE &
-Sched.
-mAP@0.5 Size (KB) mAP@0.5 Size (KB)
-- 0.93 26,292.36 0.55 22,670.62
-- 0.94 13,516.80 0.61 7,374.995
-0.92 43.84 0.69 279.97
-Impact of Core Components.
-Table 3 ablates our two main contributions: the Semantic Loss $\mathcal{L}_{\text{semantic}}$ and the FUE & Scheduling module.
-Removing the FUE & Scheduler still allows the encoder’s $\mathcal{L}_{\text{semantic}}$ to create sparse feature maps.
-However, the entire maps are transmitted, leading to redundancy and a large payload.
-Without $\mathcal{L}_{\text{semantic}}$, the FUE & Scheduler deal with dense and semantically-overloaded features, making it hard to select due to noisy utility estimates, still resulting in a large payload.
-Only with the presence of both components, the transmission size is greatly reduced while attaining high CP accuracy.
-Effect of Redundancy.
-The system’s transmission model is based on Theorem (1), demonstrating that selecting features from more than the Top-1 agent per cell is suboptimal.
-Fig.
-8 validates this empirically.
-Comparing our Top-1 Agent policy to a Top-2 Agents policy, results show that Top-2 is worse at nearly every bandwidth point, indicating that transmitting redundant features wastes bandwidth and degrades perception accuracy.
-Although adding more data seems beneficial, this highlights a limitation in current fusion networks.
-Sec.
-5 discusses this phenomenon and its implications further.
-5 Discussion
-On the ‘Cost’ of Redundancy.
-A critical, counterintuitive finding from our evaluation is that transmitting from the Top-2 agents per cell yields worse performance than transmitting the Top-1 agent alone, as shown in Fig.
-8.
-This ‘more is worse’ outcome stems from a fundamental limitation presented in current SOTA CP fusion algorithms, such as element-wise max-pooling or dot-product attention [xu2022opv2v, hu2022where2comm, f_cooper].
-The fused feature is a weighted sum of the received features, with the informative activations in the fused feature upper-bounded by the maximum signal across the individual features.
-Therefore, unlike human vision, these methods struggle to synergetically fuse redundant information.
-If the second-best agent’s feature is noisier or semantically different (e.g.
-conflicting channel activations), it can ‘dilute’ the clean, high-utility representation from the Top-1 agent.
-This finding corroborates the design philosophy of JigsawComm, which selects only the highest-quality, non-redundant data, resulting in smaller overhead and higher accuracy.
-(a) OPV2V
-(b) DAIR-V2X
-Figure 8: Ablation on number of agents sharing per cell. Including the agent with the second-highest utility does not improve the utility-bandwidth tradeoff, which corroborates Theorem (1).
-Why Not a Naive Communication Loss?
-A straightforward approach to reduce data size might be to add a simple communication loss directly on the transmission masks $\{\mathbf{M}\}_{i=1}^{N}$, penalizing the total number of transmitted bits.
-We found this approach to produce suboptimal performance in practice.
-Such a loss disregards the semantic importance and redundancy of the information transmitted, yet penalizes all communication.
-Our framework avoids this by training the FUE head to learn a task-aligned utility proxy, ensuring that selection decisions are driven by semantic-importance and redundancy.
-Relation to Semantic Communications.
-JigsawComm is related to semantic/goal-oriented communications [strinati2021_6g, binucci2022adaptive, pezone2022goal], since they share the high-level goal of task-oriented data compression.
-However, existing work in semantic communications targets different applications [strinati2021_6g], considers only single sensor perception [yuan2023scalable, binucci2022adaptive, pezone2022goal], or neglects sensor correlations [sheng2024semantic].
-Applying semantic communications to CP involves unique challenges due to its collaborative and multi-source input nature, where inter-agent redundancy prevails due to overlapping sensory data.
-Essentially, the CP problem studied in this paper can be reformulated as a joint, multi-source, task-oriented data compression problem, which is difficult to solve without knowing a priori the exact inter-agent data correlation structure.
-JigsawComm solves this problem via a data-driven approach (with a small overhead of exchanging semantic feature utilities).
-Meanwhile, the semantic regularization and redundancy-eliminating transmission scheduling design of JigsawComm can be viewed as a way of implementing the information bottleneck (IB) principle, where unnecessary information is minimized while task-relevant information is maximized, which is essential to achieve goal-oriented communications [strinati2021_6g, pezone2022goal].
-Interestingly, the same IB principle has also been shown to be effective in reducing generalization errors in deep learning models [kawaguchi2023does].
-Here, our end-to-end trained JigsawComm learns to keep only semantically-relevant information from its sparse encoder and selects non-redundant and high-utility features for transmission, which effectively controls the IB of the CP pipeline.
-Therefore, in addition to the drastic reduction in required bandwidth for real-time communication, JigsawComm also achieves matching or improved perceptual quality compared to full-feature baselines.
-Practical Implementation.
-Our framework is designed for practical V2X implementation by leveraging its dual-channel architecture [FCC_DSRC, CV2X_ETSI_EN_303_613].
-The meta utility maps, though scaling linearly with agent count, can be transmitted over the control channel, incurring negligible overhead (<<0.03 KB per agent under 4-bit quantization).
-The final, sparse BEV features are then sent efficiently over the data channel, such as via the Collective Perception Message (CPM) defined by ITS [cpm].
-Note that, our scheduler outputs the transmission priorities of all agents’ selected features as a byproduct (ordered by the corresponding feature utility values), which can be easily integrated into the low-layer V2X protocols for both scheduling-based (e.g.
-C-V2X, 5G NR-V2X [CV2X_ETSI_EN_303_613, garcia2021tutorial]) and random-access V2X protocols (e.g.
-DSRC [FCC_DSRC]).
-Our scheme is compatible with both centralized (V2I) and decentralized (V2V) operating modes.
-For V2I, an edge server co-located with the RSU can perform scheduling, fusion, and inference.
-For V2V, the same set of computations is performed by each agent in a distributed manner.
-6 Conclusion and Future Work
-We present JigsawComm, a novel joint semantic feature encoding and transmission framework for CP that maximizes every transmitted bit’s contribution to the final task accuracy.
-By jointly training a semantic sparse encoder and a lightweight Feature Utility Estimator via a differentiable scheduler, JigsawComm learns to extract and select only essential, non-redundant data.
-This enables a provably optimal top-1 policy at inference that eliminates cross-agent redundancy and achieves a scalable $\mathcal{O}(1)$ communication cost.
-Our experiments on OPV2V and DAIR-V2X demonstrate SOTA accuracy with up to over $500\times$ bandwidth reduction.
-There are several promising directions for future work.
-First, as JigsawComm adopts the index-based encoding to convert sparsely encoded features to bits, we will further push the limit of JigsawComm’s bandwidth efficiency by integrating traditional lossless source coding methods such as entropy coding [storer1988data].
-Second, we will also extend JigsawComm with joint semantic source and channel coding to deal with unreliable channel conditions.
-Third, we will study the impact of different sensor modalities and feature cell granularities on the accuracy-bandwidth tradeoff, and explore novel fusion networks that can leverage, rather than suffer from, redundancy.
-Finally, we will study practical V2X-based CP system design, robustness to communication failures or packet losses, and end-to-end goal-oriented CP systems for autonomous driving.
+Appendix 0.E Robustness to Packet Loss
+In real-world V2X deployments, transmitted feature payloads may be lost due to channel fading, congestion, or collisions at the MAC layer.
+We study how JigsawComm behaves when a non-ego agent’s scheduled feature packet is dropped with probability $p_{\text{drop}}\in\{0,0.05,0.10,0.15,0.20\}$.
+Fallback mechanism.
+Under the top-1-per-cell policy, each spatial cell is assigned to exactly one agent.
+When a non-ego agent’s packet is lost, the ego vehicle falls back to its own local feature at the affected cell.
+Since the ego’s feature is always available (no transmission required), this fallback incurs zero additional communication and requires no retraining.
+Alternatively, the system can opt for Top-$K$ scheduling when the bandwidth is sufficient.
+Results.
+Table 4 reports the results on OPV2V.
+JigsawComm exhibits remarkable resilience: even at a 20% packet drop rate, AP@0.5 decreases by only 1.4 percentage points (from 0.925 to 0.911), and AP@0.7 by 1.8 points (from 0.826 to 0.808).
+The degradation is nearly linear in $p_{\text{drop}}$, confirming that there is no catastrophic failure mode.
+This resilience stems from two properties of JigsawComm’s design: (i) the top-1 scheduler assigns each cell to the single highest-utility agent, so the dropped feature is typically from a non-ego agent that only marginally outperforms the ego’s own observation at that cell (due to less occlusion or high-quality observation); and (ii) the ego’s local feature, while not the top-ranked choice, still provides a semantically meaningful representation for non-occluded objects.
+Top-$K$ as an alternative.
+An alternative strategy for handling packet loss, especially for enhancing the detection of occluded objects, is to relax the top-1 policy to a Top-$K$ policy, where the scheduler pre-selects $K$ candidate agents per cell and falls back to the next-best candidate upon packet loss.
+However, Top-2 scheduling can incur additional transmission overhead.
+Therefore, the ego-fallback mechanism is preferable: it achieves the full accuracy benefit of Top-1 scheduling when packets arrive, and gracefully falls back to the ego’s own representation otherwise—all without increasing the nominal bandwidth or requiring changes to the scheduling policy.
+Appendix 0.F Qualitative Results
+JigsawComm ERMVP Where2comm
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Refer to caption Refer to caption Refer to caption
+Figure 10: Qualitative comparison of 3D detection results on OPV2V.
+Each row shows one scene; columns correspond to JigsawComm, ERMVP [Zhang_2024_CVPR], and Where2comm [hu2022where2comm].
+Green: ground truth; Red: predictions.
+JigsawComm produces tighter bounding boxes with fewer missed detections, especially for distant or partially occluded vehicles.